### v1

# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

Abstract
Video Super-Resolution (VSR) reconstructs high-resolution videos from low-resolution inputs to restore fine details and improve visual clarity. While deep learning-based VSR methods achieve impressive results, their centralized nature raises serious privacy concerns, particularly in applications with strict privacy requirements. Federated Learning (FL) offers an alternative approach, but existing FL methods struggle with low-level vision tasks, leading to suboptimal reconstructions. To address this, we propose FedVSR [^1] Preprint. Under review., a novel, architecture-independent, and stateless FL framework for VSR. Our approach introduces a lightweight loss term that improves local optimization and guides global aggregation with minimal computational overhead. To the best of our knowledge, this is the first attempt at federated VSR. Extensive experiments show that FedVSR outperforms general FL methods by an average of 0.85 dB in PSNR, highlighting its effectiveness. The code is available at: https://github.com/alimd94/FedVSR
1 Introduction
Refer to caption (a) Frame 029, Clip 011, REDS[35]
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[31] Refer to caption (d) FedProx[23]
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[18] Refer to caption (g) FedVSR (Ours)
Figure 1: Comparison of RVRT[24] output trained with different Federated Learning algorithms. General FL methods struggle to reconstruct fine details and textures.
Video Super-Resolution (VSR) is an advanced computational process that reconstructs high-resolution (HR) video sequences from low-resolution (LR) counterparts leveraging both spatial and temporal information [27]. Unlike Single Image Super-Resolution (SISR), which relies solely on intra-frame details, VSR exploits temporal correlations across consecutive frames to enhance the quality and sharpness of the output. This technique is particularly valuable in real-world applications such as surveillance [10], video streaming [20], and medical imaging [40, 37], where improving visual fidelity and maintaining temporal coherence are critical. VSR addresses complex and diverse degradations commonly found in low-quality videos, including down-sampling [54], noise [57], blur [33], and compression artifacts [9], while ensuring the perceptual quality and consistency of the restored video. However, the data-driven and centralized nature of existing deep learning-based VSR methods introduces a potential risk of privacy breaches, particularly when handling sensitive user data that require strict confidentiality.
Federated Learning (FL) presents a compelling alternative to mitigate privacy risks associated with centralized training in Deep Learning-based VSR methods. As a privacy-preserving framework, FL allows multiple participants or edge devices to collaboratively train a shared Machine Learning model without sharing raw data, effectively addressing data security concerns [31]. Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one. This approach minimizes communication overhead and adapts to the limited computational resources of edge devices. By avoiding centralized data collection, FL ensures the protection of sensitive information, offering a practical solution for privacy-aware applications in video super-resolution and similar domains.
Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as Image Classification [18, 23, 46], Face Recognition [36, 29], and Segmentation [32, 29]. However, their application in Super-Resolution (SR) remains largely unexplored. Studies such as [34, 51] indicate that directly applying existing FL techniques to SISR results in poor-quality reconstructions with blurred textures. This limitation arises from their inability to effectively capture rich low-level information. Motivated by this observation, we propose an FL-based approach for VSR that aims to generate better quality output while ensuring data privacy.
In this paper, we propose FedVSR a novel model-agnostic framework for VSR under the FL paradigm, which treats the VSR models as a black-box, making it highly adaptable to different VSR methods and future advancements. To the best of our knowledge, this is the first attempt to address the VSR problem within a decentralized setting. Moreover, our framework does not retain previous model weights for aggregation. Instead, we introduce an effective loss term based on Discrete Wavelet Transform (DWT) which not only enhances local training efficiency with negligible computational overhead but also informs the global aggregation process, leading to improved reconstruction quality. This design enables our approach to efficiently capture essential information while ensuring strict data privacy, making it well-suited for real-world applications where both visual fidelity and confidentiality are critical. In summary, the main contributions of this paper are as follows:
  • We propose a DWT-based loss term to enhance high-frequency detail in Federated VSR. This term also serves as a reliable performance metric for weighting parameter aggregation.
  • Our approach is fully model-agnostic and stateless, requiring no architecture-specific assumptions or storage of past weights, making it practical and scalable.
  • Extensive experiments validate our method’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in a federated setting.
  • To the best of our knowledge, this is the first work addressing federated VSR, balancing privacy preservation with improved visual results.
2 Related Works
In this section, we begin by discussing the most significant contributions in general FL and VSR. We then explore studies that specifically focus on applying FL to the SR task.
2.1 Federated Learning
Federated learning (FL) is a decentralized learning paradigm that enables multiple clients to collaboratively train models while preserving data privacy. It has been widely applied across diverse domains, including healthcare, finance, and IoT, where data security and regulatory constraints are critical [1, 5, 56]. A foundational approach in this domain is FedAvg [31], which follows a four-step iterative process: (1) the server distributes a global model to all participating clients, (2) each client trains its local model using its own data, (3) the locally updated models are sent back to the central server, and (4) the server aggregates these updates to refine the global model for the next training round.
To enhance the performance of federated learning, multiple methods focus on optimizing local updates. FedProx [23] incorporates a proximal term in the local objective to ensure consistency between local and global models. SCAFFOLD [18] introduces control variates for both clients and the server. FedNova [46] employs normalized averaging to address objective inconsistencies. More recently, MOON [21] applies contrastive learning to assess similarity among different model representations.
2.2 Video Super Resolution
Early VSR methods that applied SISR techniques frame-by-frame neglected the temporal context, resulting in temporal inconsistencies, loss of detail, and undesired smoothing in the high-resolution output [28]. VSR leverages multiple temporally correlated low-resolution frames within a video sequence to generate high-resolution output. Unlike SISR, VSR must account for both spatial and temporal dependencies across frames, making it a highly complex, non-linear, multi-dimensional problem.
Recent advancements in VSR models have introduced various techniques to enhance performance. VRT [25] enables parallel processing of long video sequences using a mutual attention mechanism for feature extraction, alignment, and fusion. RVRT [24] combines parallel and recurrent approaches, employing guided deformable attention for precise clip alignment. IART [49] uses implicit resampling with local cross-attention to preserve frequency details while reducing spatial distortions. DiQP [9], the first model for 8K videos, employs Denoising Diffusion to counteract compression artifacts while maintaining robustness across varying levels.
2.3 Federated Learning on Super Resolution Task
For low-level vision tasks, FedMRI [13] introduces a federated learning framework for MRI reconstruction with a shared encoder for generalized features and client-specific decoders for personalization. FedPR [14] improves image quality using pre-trained models and prompt learning with minimal trainable parameters. p²FedSR [50] enhances privacy-preserving SR by integrating client-specific degradation priors and a selective model aggregation strategy for personalized FL. FedSR [51] employs a detail-assisted contrastive loss for feature alignment at the client level and a hierarchical aggregation policy at the server level with historical global model, optimizing layer-wise model integration.
3 Preliminaries
3.1 VSR Problem Definition
VSR enhances video resolution by using multiple low-resolution frames, leveraging temporal information to exploit inter-frame dependencies, unlike SISR.
Problem Formulation
In addition to the widely used RGB color space, other representations such as the YUV and YCbCr color spaces are frequently employed in VSR applications. Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in a given LR video sequence, while $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ represents its corresponding high-resolution (HR) counterpart, where $s$ is the scaling factor (e.g., $s=2,4,8$). Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius. The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as: {scriptsize }
$I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha})$ (1)
where $\phi(\cdot;\cdot)$ is the degradation function parameterized by $\theta_{\alpha}$, incorporating various degradation factors such as noise, motion blur, and down-sampling effects.
A more detailed expression for the degradation process can be written as: {scriptsize }
$I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j}$ (2)
where $D$ represents the down-sampling operation, $B$ denotes the blurring effect, $E_{i\to j}$ is a warping function that accounts for motion between frames $\hat{I}_{i}$ and $\hat{I}_{j}$, and $n_{j}$ represents the image noise. The goal of a VSR algorithm is to reconstruct an HR video sequence from degraded LR frames in a manner that closely resembles the ground truth video. This makes this task an ill-posed problem because there exist infinitely many possible HR images $\hat{I}_{i}$ that can produce the same LR image $I_{i}$ after degrading, meaning the degradation operator is highly non-injective [44]. While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence. The inverse process of degradation, which represents the super-resolution task, can be expressed as: {scriptsize }
$\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta})$ (3)
where $\tilde{I}_{i}$ is the estimated HR frame corresponding to the ground truth $\hat{I}_{i}$, and $\theta_{\beta}$ represents the model parameters involved in the reconstruction process [28].
3.2 FL Problem Definition
FL enables multiple clients to train a shared model collaboratively without sharing raw data, enhancing privacy and reducing communication overhead through local computations and periodic model aggregation.
Problem Formulation
FL aims to minimize the following objective function across participating clients: {scriptsize }
$\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w})$ (4)
where:
  • $K$ is the total number of clients,
  • $p_{k}$ is the relative weight of client $k$ (typically based on its dataset size),
  • $F_{k}(\mathbf{w})$ is the local loss function for client $k$,
  • $\mathbf{w}$ represents the global model parameters.
Each client updates the global model locally: {scriptsize }
$\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t})$ (5)
where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme: {scriptsize }
$\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}$ (6)
Not all clients participate equally in every training round due to data heterogeneity, resource constraints, network instability, or other limitations [53]. Client averaging based on their contributions stabilizes training, reduces variance, and improves generalization [4].
However, weighting clients solely by dataset size, as in traditional FedAvg[31], does not fully capture their impact on model performance [39]. A client’s importance extends beyond data quantity and is also influenced by the quality of its model updates and the loss it generates during training. Based on [41], we implement an approach where clients with lower loss are typically better trained and can be assigned higher weights during averaging. This ensures improved overall performance by leveraging the robustness properties analyzed in the paper. It is worth noting that some works instead prioritize clients with higher loss to increase fairness [22]. This approach, which contrasts with a focus on final performance, is more common in the context of personalized federated learning.
By incorporating loss-based importance into the aggregation process, federated learning can achieve more reliable and efficient training. This adaptive weighting ensures that client updates more accurately reflect their significance in improving the global model, leading to a more stable and equitable learning process [15].
4 Methodology
VSR remains a challenging problem due to its ill-posed nature. Severe degradations amplify artifacts, making it difficult to balance detail enhancement and noise suppression. To improve generalizability, diverse training data and large batch sizes are needed, but this significantly increases computational costs [6]. FL adds further complexity to VSR due to real-world constraints [12] and the nature of the task. In heterogeneous environments, differences in client data distributions lead to biased updates, poor generalization [3], and inefficient model merging. Gradient mismatches [48] and parameter inconsistencies slow convergence [8]. Most FL algorithms are designed for image classification which can be considered as a well-posed problem [16]. However, VSR is an ill-posed task, requiring a further fundamental assessment of existing methods to ensure effective adaptation. For instance, FedProx assumes a certain degree of similarity across client data distributions, but in real-world applications, these distributions can vary significantly. When this assumption fails, models experience drift, leading to inconsistent local updates and poor generalization [11]. Similarly, SCAFFOLD assumes unbiased client updates, yet in large-scale deployments, heterogeneous data causes significant divergence, slowing convergence. Moreover, its need for additional state information per client makes it impractical for large-scale FL systems due to communication and storage constraints [8]. Meanwhile, Other methods such as FedDisco[52] and MOON rely on class label distributions or logits, making them inherently task-specific rather than generalizable.This work proposes a solution to enhance the overall objective of VSR models while improving aggregation to mitigate these challenges.
4.1 Enhancing Loss Function for Federated VSR
Selecting a proper loss function as objective for VSR model is crucial and determine the overall output’s performance [45, 42] and as depicted in Fig.1. It is evident that existing FL algorithms has problems in capturing and generating details in the output. Since we aim to avoid adding computational overhead to the client by using adversarial or perceptual loss, and because our objective is to keep the model untouched, ensuring FedVSR remains model-agnostic, we introduce an additional term to the overall loss of model improve performance.
4.1.1 DWT for High-Frequency Preservation
One common failure in training VSR models is producing overly smooth outputs, meaning the trained model fails to capture high-frequency information. As a result, details and textures are missing from the output. Since the models we used are among the top-performing ones in a centralized setting, this suggests that the failure to restore high-frequency details is specifically related to training under the FL paradigm. To alleviate this, we leverage the Discrete Wavelet Transform (DWT), which has proven effective in super-resolution tasks [17, 55].
DWT is a multi-resolution signal decomposition technique that represents a signal in terms of approximation and detail coefficients at different scales. For a 2D signal (such as an image or video frame), DWT decomposes the signal into sub-bands of different frequencies and orientations such as:
  • Approximation Coefficients $F_{i}^{LL}$ (low-pass filtered in both dimensions).
  • Horizontal Detail Coefficients $F_{i}^{LH}$ (high-pass in vertical, low-pass in horizontal).
  • Vertical Detail Coefficients $F_{i}^{HL}$ (low-pass in vertical, high-pass in horizontal).
  • Diagonal Detail Coefficients $F_{i}^{HH}$ (high-pass in both dimensions).
One of the simplest wavelets used in DWT is Haar wavelet, which consists of a step function. It is computationally efficient and commonly used in image and video processing tasks. In this method each frame’s pixel values are convolved with:
  • Low-pass filter $h=\frac{1}{\sqrt{2}}[1,1]$ to extract approximation coefficients.
  • High-pass filter $g=\frac{1}{\sqrt{2}}[1,-1]$ to extract detail coefficients.
This process is repeated independently for each frame in the sequence.
The coefficient in those four sub-bands are calculated as: {scriptsize }
$F_{i}^{LL}=(F_{i}*h^{T})*h,\quad F_{i}^{LH}=(F_{i}*h^{T})*g$ (7)
{scriptsize }
$F_{i}^{HL}=(F_{i}*g^{T})*h,\quad F_{i}^{HH}=(F_{i}*g^{T})*g$ (8)
The low-frequency component $F_{i}^{LL}$, contains the content and color information of the input frame, while the three high-frequency sub-bands capture detailed information about global structures and textures [38]. Therefore, for our loss helper term, we consider only the high-frequency sub-bands and ignore the low-frequency ones.
The loss is then computed pixel-wise in each sub-band between the predicted $\bar{F}$ and ground truth $\hat{F}$ frame at time step $i$:
{scriptsize }
$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH},\quad L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL},\quad L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (9)
Then with the Charbonnier loss [7] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ : {scriptsize }
$\mathcal{L}_{\text{HiFr}}=\sum_{i}\sum_{s\in\{LH,HL,HH\}}\sqrt{{L_{i}^{s}}^{2}+\epsilon^{2}}$ (10)
Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows: {scriptsize }
$\mathcal{L}_{\text{total}}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$ (11)
As our experiments demonstrate (see Section 5.3), the optimal weight values are unbounded, meaning they are not constrained by the relationship $\lambda_{\text{VSR}}=1-\lambda_{\text{HiFr}}$ and can be independently adjusted for better performance.
4.2 Loss-Aware Aggregation for Federated VSR
To further enhance the model’s performance, we employ a greedy aggregation approach with a decay rate, inspired by MKL-SGD [41]. MKL-SGD improves robustness by selecting samples with the lowest loss during training, thereby mitigating the impact of outliers. Similarly, in our method, clients accumulate their loss values throughout local training and compute an average loss based on the number of training steps. When sending their updated model weights to the server, they also transmit this average loss. At the server, we perform inverse weighted averaging, where each client’s weight is computed as $I_{i}=\frac{1}{\mathcal{L}_{i}}$ and adjusted using an adaptive step: $I^{\prime}_{i}=I_{i}^{\alpha}$ where $\alpha$ controls the strength of loss-based weighting. To regulate this process, we introduce a decay rate that gradually reduces the effect of loss-based weighting over time, following:$\alpha=\alpha^{1-\frac{t}{T}}$ where $t$ is the current round and $T$ is the total number of rounds. This ensures that early training benefits from loss-based prioritization, while later rounds transition smoothly toward standard averaging, balancing robustness and generalization. The overall algorithm is summarized in Algorithm 1.
Algorithm 1 Our Method
1:Number of communication rounds $T$, Number of clients $N$, Learning rate $\eta$, Adaptive step $\alpha$, Decay rate $\gamma$
2:Global model weights $W_{T}$
3:Server Initialization Initialize global model $W_{0}$
4:for $t=1,\dots,T$ do
5:    Server selects a subset of clients randomly $S_{t}\subseteq\{1,\dots,N\}$
6:    for each client $i\in S_{t}$ in parallel do
7:        Client updates local model:
8:         Retrieve latest global model $W_{t}$
9:         Perform local training for $E$ epochs with learning rate $\eta$
10:        Compute Local Loss:
11:         Compute model-specific loss $\mathcal{L}_{\text{VSR}}$
12:         Compute $\mathcal{L}_{\text{DWT}}$ loss $\mathcal{L}_{\text{DWT}}$
13:         Compute total loss
$\mathcal{L}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$
14:        Client updates local model:
$\mathbf{w}_{i}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{i}(\mathbf{w}^{t})$
15:         Return updated weights $W_{i}$ and loss value $\mathcal{L}_{i}$ to server
16:    end for
17:    Server Aggregation: (Performed at the Server)
18:     Compute inverse loss weighting $I_{i}=\frac{1}{\mathcal{L}_{i}}$
19:     Apply adaptive step $I^{\prime}_{i}=I_{i}^{\alpha}$
20:     Normalize inverse loss weights $w_{i}=\frac{I^{\prime}_{i}}{\sum_{j\in S_{t}}I^{\prime}_{j}}$
21:     Compute weighted model update $W_{i}^{\text{scaled}}=w_{i}\cdot W_{i}$
22:     Aggregate global model $W_{t+1}=\sum_{i\in S_{t}}W_{i}^{\text{scaled}}$
23:    Apply Decay Rate to Adaptive Step $\alpha=\alpha^{1-\frac{t}{T}}$
24:end for
25:return Global model $W_{T}$
5 Experiments
Model Frames, Params(M) Algorithm Less Heterogeneous More Heterogeneous
REDS[35] Vid4[26] REDS[35] Vid4[26]
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
VRT[25] 16, 35.6 FedAvg[31] 29.71 0.8522 24.88 0.7680 28.85 0.8282 24.60 0.7516
FedProx[23] 28.82 0.8281 24.18 0.7341 28.64 0.8236 23.99 0.7256
SCAFFOLD[18] 28.59 0.8181 23.99 0.7182 28.17 0.8074 24.04 0.7208
FedVSR (Ours) 29.86 0.8562 24.92 0.7682 29.03 0.8339 24.73 0.7584
RVRT[24] 30, 10.8 FedAvg[31] 31.30 0.8877 25.96 0.8156 29.89 0.8578 25.67 0.8040
FedProx[23] 30.09 0.8618 25.20 0.7775 29.59 0.8526 25.02 0.7730
SCAFFOLD[18] 29.63 0.7967 24.79 0.7317 26.71 0.7142 23.08 0.6441
FedVSR (Ours) 31.49 0.8908 26.05 0.8191 30.34 0.8678 25.90 0.8048
IART[49] 16, 13.4 FedAvg[31] 30.56 0.8729 25.43 0.7949 29.52 0.8493 25.04 0.7774
FedProx[23] 29.63 0.8500 24.72 0.7599 29.30 0.8444 24.42 0.7505
SCAFFOLD[18] 30.29 0.8678 25.18 0.7857 29.26 0.8413 24.85 0.7690
FedVSR (Ours) 30.88 0.8812 25.54 0.8012 29.70 0.8518 25.26 0.7850
Table 1: PSNR and SSIM comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
5.1 setup
Dataset:
Following [25, 24, 49], one of the most commonly used datasets for training VSR is REDS [35]. It contains 240, 30, and 30 sequences for training, validation, and testing, respectively, with each sequence consisting of 100 frames at a resolution of 720 × 1280. The corresponding low-resolution (LR) frames are generated through bicubic downsampling from the high-resolution (HR) images. This dataset features a variety of natural and handmade objects, locations, and diverse content, making it heterogeneous in decentralized settings [35]. To further evaluate FedVSR and assess its robustness in more realistic, highly heterogeneous scenarios, we conducted additional experiments using Kinetics-400 [19], a dataset primarily used for human action recognition. It includes 400 action classes with 306,245 videos, each approximately 10 seconds long, sourced from unique YouTube videos. Since Kinetics-400 contains videos of varying resolutions, we selected only those with a 720 × 1280 resolution to ensure a fair comparison with REDS. From the 400 classes, we randomly chose 40, and for each class, selected six videos, extracting 100-frame clips to match REDS’ format. Kinetics-400 is particularly well-suited for testing FedVSR under more extreme heterogeneity because, unlike professionally filmed datasets, its videos often exhibit camera shake, lighting variations, shadows, and background clutter [19]. This makes it a challenging but realistic benchmark for assessing the model’s adaptability in diverse and less controlled environments. For consistency with our experimental setup, we refer to REDS as the ”less heterogeneous” dataset and Kinetics-400 as the ”more heterogeneous” dataset, given the increased variability in content, motion, and environmental conditions in the latter.
Implementation Details:
To maintain our objective of developing a model-agnostic framework, we did not alter the official parameters of the baseline VSR models. Instead, we used their recommended settings for training. We created 40 clients, each assigned six videos in both experimental setups. In the more heterogeneous setup, each client received videos from a single class; for example, one client had only ”archery” videos, while another had only ”making pizza” videos. In contrast, in the less heterogeneous setup, the six videos were randomly assigned to each client. We set the total number of global rounds to 100, with each client performing one local round (epoch) per global round. In each round, we randomly selected four clients to perform training. After completing their local training, these clients sent their updated weights to the server for aggregation. All experiments were conducted using bi-cubic (BI) degradation.
Baselines:
For our baseline, we use a combination of FedAvg[31], FedProx[23], and SCAFFOLD[18] from the FL side, along with VRT[25], RVRT[24], and IART[49] from VSR.
Evaluation Metrics:
The quality of the reconstructed video is typically evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM).
Figure 2: PSNR values across different rounds for various test sets under different settings for VRT [25], RVRT [24], and IART [49].
Reproducibility:
To ensure our experiments are reproducible, we fixed the random seed, which doesn’t affect client selection, model initialization, etc. We used four NVIDIA A6000 GPUs for training, with each model under a federated learning algorithm taking an average 46 hours to train. For the baselines alone, this consumed 36 days (2 experiments × 9 models+algorithms × 2 days per run).
5.2 Results and Comparison
Quantitative Results:
As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets. On average, FedVSR surpasses existing FL methods by a margin of 0.89 dB in PSNR and 0.028 in SSIM in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.80 dB in PSNR and 0.027 in SSIM, demonstrating its robustness under challenging conditions. Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin. In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods. When comparing FedVSR only against the second-best FL algorithm, FedAvg, our approach achieves an average PSNR margin of 0.22 dB in less heterogeneous settings and 0.27 dB in more heterogeneous settings. Similarly, the SSIM improvement over FedAvg is 0.005 in less heterogeneous cases and 0.006 in more heterogeneous cases. The smallest improvement relative to FedAvg is observed for VRT in PSNR and IART in SSIM, while the highest improvement is consistently achieved by RVRT, particularly in more heterogeneous settings, further validating FedVSR’s ability to handle real-world federated VSR challenges effectively. The findings emphasize the importance of high-frequency-aware optimization in VSR FL-based training.
The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results. As shown in Fig.2, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FedAvg, FedProx, and SCAFFOLD, particularly in later training rounds. This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions. The study suggests that FedProx’s limiting points may remain far from the true minimizer of the global empirical risk [43], which could explain its suboptimal performance in our experiments. Although FedProx may still generalize well by aggregating local data effectively in certain conditions, its weaker convergence properties become apparent in more heterogeneous scenarios. We also observe that FedProx is highly sensitive to the $\mu$ value during training. The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases. However, we notice that for any $\mu$ greater than $31e-$ 3, the VSR model struggles significantly with learning. This suggests that the relationship is more complex and requires further investigation and analysis.
Similarly, our findings regarding SCAFFOLD’s inefficacy in highly heterogeneous settings corroborate previous research that highlights its susceptibility to instability under heterogeneous data [2]. The significant performance drops and stagnation observed in our experiments, particularly with the RVRT model, align with the reported issues of SCAFFOLD transmitting nearly twice the data of other algorithms due to its reliance on control variates. This added communication burden does not necessarily translate to better convergence, as our results show that SCAFFOLD struggles to optimize effectively in diverse settings. Prior work [2] also notes that SCAFFOLD is highly prone to sudden performance breakdowns when gradient clipping is not employed, which is reflected in our results where its PSNR fluctuates or plateaus early in highly heterogeneous scenarios. The increased variance in SCAFFOLD’s performance under skewed data distributions further supports our findings that it fails to adapt smoothly across different client datasets, reinforcing its limitations in federated VSR tasks.
Overall, these existing studies provide strong theoretical support for our experimental results. In contrast, FedVSR’s convergence, stability, and robustness across varying levels of data heterogeneity highlight its effectiveness in overcoming these limitations.
Qualitative Results:
As showen in Fig. 1 clearly demonstrate the limitations of FedAvg, FedProx, and SCAFFOLD in accurately reconstructing fine details. While these methods reduce some pixelation from the LQ input, they all fail to recover the textures and sharp edges present in the ground-truth. FedAvg produces overly smooth results, blurring important structural details such as the wooden lattice and frame edges. FedProx exhibits similar blurring, struggling to restore high-frequency textures. SCAFFOLD, not only fails to produce a sharp reconstruction but also introduces noticeable artifacts, distorting the texture of the wooden lattice and making the output appear unnatural. In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches. Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 7.
5.3 Ablation Study
Evaluating $\mathcal{L}_{\text{HiFr}}$ Loss and Adaptive Aggregation
To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments. Using FedAvg as a baseline, we added each component separately to evaluate their individual effects. To assess the impact of the $\mathcal{L}_{\text{HiFr}}$ term, we applied it to FedAvg and observed the changes (F-$\mathcal{L}_{\text{HiFr}}$). Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds. We then compared this approach to an adaptive version that incorporates a decay mechanism (F-ALA). As shown in Table 2, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases. Additionally, both methods individually surpass FedAvg [31]. However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics. These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights.
Model Setting Less Heterogeneous More Heterogeneous
REDS Vid4 REDS Vid4
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 24.91 0.7681 28.98 0.8322 24.70 0.7577
F-ALA 29.75 0.8531 24.89 0.7681 28.90 0.8313 24.64 0.7559
F-GLA 29.70 0.8523 24.86 0.7670 28.83 0.8279 24.55 0.7500
RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 26.02 0.8178 30.24 0.8617 25.84 0.8042
F-ALA 31.38 0.8880 26.00 0.8171 30.19 0.8601 25.79 0.8041
F-GLA 31.31 0.8879 25.95 0.8155 29.88 0.8576 25.60 0.8032
IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 25.46 0.8001 29.64 0.8508 25.21 0.7837
F-ALA 30.70 0.8789 25.44 0.7998 29.60 0.8500 25.18 0.7831
F-GLA 30.57 0.8730 25.42 0.7951 29.50 0.8482 24.99 0.7761
Table 2: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR and SSIM under different heterogeneity settings.
Optimal $\lambda$s
Using a grid search, we explored different values for $\lambda_{\text{VSR}}$ and $\lambda_{\text{HiFr}}$ to determine the optimal ones. We varied each parameter from 1 to 0.1 in steps of 0.1 under both bounded and unbounded conditions. Due to the high computational cost of the experiments, we trained the VSR models for only 20 rounds and selected the best-performing configurations. Our results show that for VRT[25], the optimal values for both parameters are 1 For RVRT[24], the best values are 1 and 0.5, while for IART[49], the optimal values are 0.5 and 0.5, respectively.
Empirical Fisher for Estimating Weights Importance
Inspired by [30], we explored estimating the importance of each weight in client models to enhance aggregation using Empirical Fisher (EF) Information. EF approximates the Fisher information matrix by leveraging gradients collected during backpropagation after training. While this method led to slight performance improvements, it proved computationally too expensive for clients, making it impractical for our setting. A more efficient variant of EF was introduced in [47]; however, as it is tailored for classification tasks, we leave its adaptation to VSR for future work.
6 Conclusion
In this work, we introduced FedVSR, the first FL framework specifically designed for VSR, addressing privacy concerns while preserving high-quality performance. Unlike existing FL approaches that struggle with low-level vision tasks, our model-agnostic and stateless design ensures adaptability across various VSR architectures without prior model knowledge or historical weight storage. By incorporating a lightweight loss term, we effectively guide local updates and enhance global aggregation, resulting in notable performance improvements over general FL baselines. Our method functions as a plug-and-play solution, and extensive experiments confirm its superiority over existing FL approaches, demonstrating its potential for real-world, privacy-sensitive applications. Furthermore, our study highlights key challenges in applying FL to low-level vision tasks, offering valuable insights for future research in this under-explored domain. As the first FL-based approach for VSR, this work establishes a strong baseline and provides a clear research path for future advancements. Despite promising results, a gap remains between FL-based VSR and centralized approaches. Future research could refine the aggregation strategy with a more intelligent, task-specific approach tailored for low-level vision. Additionally, our loss-guided optimization could be further enhanced to improve stability and better address existing challenges in FL-based VSR.
Supplementary Material
7 Further Analysis: Performance Metrics and Visual Comparisons
This section provides additional performance evaluations for our proposed method. Figures 3, 4, and 5 illustrate the PSNR and SSIM trends across different rounds for multiple test sets under various configurations of VRT [25], RVRT [24], and IART [49]. These results further highlight the effectiveness of our approach in enhancing the performance across different models.
Figures 6, 7, and 8 provide qualitative visual comparisons on various test frames. Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR). The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
Figure 3: PSNR and SSIM values across different rounds for various test sets under different settings for VRT[25].
Figure 4: PSNR and SSIM values across different rounds for various test sets under different settings for RVRT[24].
Figure 5: PSNR and SSIM values across different rounds for various test sets under different settings for IART[49].
Figure 6: Comparison of RVRT output trained with different Federated Learning algorithms under more heterogeneous setting
Figure 7: Comparison of VRT output trained with different Federated Learning algorithms under more heterogeneous setting
Figure 8: Comparison of IART output trained with different Federated Learning algorithms under less heterogeneous setting

### v2

@@
 Abstract
-Video Super-Resolution (VSR) reconstructs high-resolution videos from low-resolution inputs to restore fine details and improve visual clarity.
-While deep learning-based VSR methods achieve impressive results, their centralized nature raises serious privacy concerns, particularly in applications with strict privacy requirements.
-Federated Learning (FL) offers an alternative approach, but existing FL methods struggle with low-level vision tasks, leading to suboptimal reconstructions.
-To address this, we propose FedVSR [^1] Preprint.
-Under review., a novel, architecture-independent, and stateless FL framework for VSR.
-Our approach introduces a lightweight loss term that improves local optimization and guides global aggregation with minimal computational overhead.
-To the best of our knowledge, this is the first attempt at federated VSR.
-Extensive experiments show that FedVSR outperforms general FL methods by an average of 0.85 dB in PSNR, highlighting its effectiveness.
+Video super-resolution aims to enhance low-resolution videos by leveraging both spatial and temporal information.
+While deep learning has led to impressive progress, it typically requires centralized data, which raises privacy concerns.
+Federated learning offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs.
+To address this, we introduce FedVSR, the first FL framework specifically designed for VSR.
+It is model-agnostic and stateless, and introduces a lightweight loss function based on the DWT to better preserve high-frequency details during local training.
+Additionally, a loss-aware aggregation strategy combines both DWT-based and task-specific losses to guide global updates effectively.
+Extensive experiments across multiple VSR models and datasets demonstrate that FedVSR consistently outperforms existing FL methods, achieving up to 0.82 dB higher PSNR, 0.0327 higher SSIM, and 0.0251 lower LPIPS.
+These results underscore FedVSR’s ability to bridge the gap between privacy and performance, setting a new benchmark for federated learning in low-level vision tasks.
 The code is available at: https://github.com/alimd94/FedVSR
@@
 Refer to caption (a) Frame 029, Clip 011, REDS[40]
-Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[31] Refer to caption (d) FedProx[23]
-Refer to caption (e) GT Refer to caption (f) SCAFFOLD[18] Refer to caption (g) FedVSR (Ours)
-Figure 1: Comparison of RVRT[24] output trained with different Federated Learning algorithms. General FL methods struggle to reconstruct fine details and textures.
-Video Super-Resolution (VSR) is an advanced computational process that reconstructs high-resolution (HR) video sequences from low-resolution (LR) counterparts leveraging both spatial and temporal information [27].
-Unlike Single Image Super-Resolution (SISR), which relies solely on intra-frame details, VSR exploits temporal correlations across consecutive frames to enhance the quality and sharpness of the output.
-This technique is particularly valuable in real-world applications such as surveillance [10], video streaming [20], and medical imaging [40, 37], where improving visual fidelity and maintaining temporal coherence are critical.
-VSR addresses complex and diverse degradations commonly found in low-quality videos, including down-sampling [54], noise [57], blur [33], and compression artifacts [9], while ensuring the perceptual quality and consistency of the restored video.
-However, the data-driven and centralized nature of existing deep learning-based VSR methods introduces a potential risk of privacy breaches, particularly when handling sensitive user data that require strict confidentiality.
-Federated Learning (FL) presents a compelling alternative to mitigate privacy risks associated with centralized training in Deep Learning-based VSR methods.
-As a privacy-preserving framework, FL allows multiple participants or edge devices to collaboratively train a shared Machine Learning model without sharing raw data, effectively addressing data security concerns [31].
+Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 30.41, 0.8600, 0.1777 Refer to caption (d) FedProx[27] 30.07, 0.8531, 0.1851
+Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 26.78, 0.7148, 0.2516 Refer to caption (g) FedVSR (Ours) 32.00, 0.8889, 0.1508
+Figure 1: Comparison of RVRT[29] output trained with different federated learning algorithms.
+General FL methods struggle to reconstruct fine details and textures.
+PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
+Video super-resolution (VSR) is an advanced computational process that reconstructs high-resolution (HR) video sequences from low-resolution (LR) counterparts, leveraging both spatial and temporal information [32].
+Unlike single-image super-resolution (SISR), which relies solely on intra-frame details, VSR exploits temporal correlations across consecutive frames to enhance the quality and sharpness of the output.
+This capability is crucial in real-world applications such as surveillance [12], video streaming [23], and medical imaging [45, 42], where visual fidelity and temporal coherence are mission-critical.
+In practice, VSR must also cope with degradations such as down-sampling [60], noise [63], blur [38], and compression artifacts [11].
+However, modern deep-learning pipelines address these degradations with data-hungry networks trained on centrally collected data, but centralization itself raises privacy, regulatory, and bandwidth concerns, especially when the raw videos contain sensitive or personally identifiable information.
+Federated learning (FL) provides a principled remedy to these concerns [36].
+In many VSR scenarios, such as smartphones, surveillance, health-related domains, and remote edge sensors, videos are captured on devices with limited bandwidth and strict privacy constraints.
+Uploading hours of footage can consume tens of gigabytes of storage; for instance, even a short 10-second 8K video at 30 fps in raw format can exceed 10 GB in size [2].
+Additionally, uploading such data may violate policies and laws such as GDPR or HIPAA [19].
 Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one.
-This approach minimizes communication overhead and adapts to the limited computational resources of edge devices.
-By avoiding centralized data collection, FL ensures the protection of sensitive information, offering a practical solution for privacy-aware applications in video super-resolution and similar domains.
-Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as Image Classification [18, 23, 46], Face Recognition [36, 29], and Segmentation [32, 29].
+This approach minimizes communication overhead and accommodates the limited computational resources of edge devices.
+These combined benefits make FL a logical and effective foundation for VSR systems that face such concerns.
+Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [21, 27, 51], facial recognition [41, 34], and segmentation [37, 34].
 However, their application in super-resolution (SR) remains largely unexplored.
-Studies such as [34, 51] indicate that directly applying existing FL techniques to SISR results in poor-quality reconstructions with blurred textures.
-This limitation arises from their inability to effectively capture rich low-level information.
-Motivated by this observation, we propose an FL-based approach for VSR that aims to generate better quality output while ensuring data privacy.
-In this paper, we propose FedVSR a novel model-agnostic framework for VSR under the FL paradigm, which treats the VSR models as a black-box, making it highly adaptable to different VSR methods and future advancements.
-To the best of our knowledge, this is the first attempt to address the VSR problem within a decentralized setting.
-Moreover, our framework does not retain previous model weights for aggregation.
-Instead, we introduce an effective loss term based on Discrete Wavelet Transform (DWT) which not only enhances local training efficiency with negligible computational overhead but also informs the global aggregation process, leading to improved reconstruction quality.
-This design enables our approach to efficiently capture essential information while ensuring strict data privacy, making it well-suited for real-world applications where both visual fidelity and confidentiality are critical.
+Existing studies [39, 56] found that directly applying existing general FL techniques to SR tasks results in poor-quality reconstructions with blurred textures, mainly due to their inability to effectively capture rich low-level information such as texture, edges and motion.
+Motivated by these limitations, we propose FedVSR, a novel model-agnostic framework for VSR under FL.
+FedVSR treats VSR models as black boxes, allowing it to work with a variety of architectures without requiring architectural changes or access to intermediate layers.
+To improve performance under FL constraints, we introduce a lightweight loss term based on the Discrete Wavelet Transform (DWT), which emphasizes high-frequency details that are often lost during decentralized training.
+While high-frequency preservation is valuable for all SR tasks, its importance is amplified in VSR due to the added complexity of maintaining temporal coherence across frames.
+In federated settings, data heterogeneity and limited communication intensify this challenge.
+FedVSR addresses these issues without maintaining state between rounds, making it both scalable and practical for privacy-sensitive, resource-constrained environments.
 In summary, the main contributions of this paper are as follows:
-  • We propose a DWT-based loss term to enhance high-frequency detail in Federated VSR. This term also serves as a reliable performance metric for weighting parameter aggregation.
-  • Our approach is fully model-agnostic and stateless, requiring no architecture-specific assumptions or storage of past weights, making it practical and scalable.
-  • Extensive experiments validate our method’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in a federated setting.
-  • To the best of our knowledge, this is the first work addressing federated VSR, balancing privacy preservation with improved visual results.
+  • We design a frequency-aware loss using DWT to sharpen fine details during client updates, and introduce a complementary aggregation method to weight model contributions more effectively.
+To the best of our knowledge, this is the first algorithm specifically designed to address VSR in the FL setting, effectively bridging the gap between privacy preservation and reconstruction quality.
+  • Our approach makes no architecture-specific assumptions and does not require storing past weights, making it stateless and model-agnostic.
+This design reduces memory usage, avoids privacy risks, and eliminates synchronization overhead in federated learning on edge devices.
+  • Extensive experiments validate FedVSR’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
 2 Related Works
@@
 Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius.
-The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as: {scriptsize }
+The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as:
 $I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha})$ (1)
 where $\phi(\cdot;\cdot)$ is the degradation function parameterized by $\theta_{\alpha}$, incorporating various degradation factors such as noise, motion blur, and down-sampling effects.
-A more detailed expression for the degradation process can be written as: {scriptsize }
+A more detailed expression for the degradation process can be written as:
 $I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j}$ (2)
@@
 While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence.
-The inverse process of degradation, which represents the super-resolution task, can be expressed as: {scriptsize }
+The inverse process of degradation, which represents the super-resolution task, can be expressed as:
 $\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta})$ (3)
@@
 Problem Formulation
-FL aims to minimize the following objective function across participating clients: {scriptsize }
+FL aims to minimize the following objective function across participating clients:
 $\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w})$ (4)
@@
   • $\mathbf{w}$ represents the global model parameters.
-Each client updates the global model locally: {scriptsize }
+Each client updates the global model locally:
 $\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t})$ (5)
-where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme: {scriptsize }
+where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme:
 $\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}$ (6)
@@
 This process is repeated independently for each frame in the sequence.
-The coefficient in those four sub-bands are calculated as: {scriptsize }
+The coefficient in those four sub-bands are calculated as:
 $F_{i}^{LL}=(F_{i}*h^{T})*h,\quad F_{i}^{LH}=(F_{i}*h^{T})*g$ (7)
-{scriptsize }
 $F_{i}^{HL}=(F_{i}*g^{T})*h,\quad F_{i}^{HH}=(F_{i}*g^{T})*g$ (8)
@@
 The loss is then computed pixel-wise in each sub-band between the predicted $\bar{F}$ and ground truth $\hat{F}$ frame at time step $i$:
-{scriptsize }
-$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH},\quad L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL},\quad L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (9)
-Then with the Charbonnier loss [7] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ : {scriptsize }
+$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH}$ (9)
+$L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL}$ (10)
+$L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (11)
+Then with the Charbonnier loss [9] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ :
 $\mathcal{L}_{\text{HiFr}}=\sum_{i}\sum_{s\in\{LH,HL,HH\}}\sqrt{{L_{i}^{s}}^{2}+\epsilon^{2}}$ (12)
-Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows: {scriptsize }
+Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows:
 $\mathcal{L}_{\text{total}}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$ (13)
@@
 5 Experiments
-Model Frames, Params(M) Algorithm Less Heterogeneous More Heterogeneous
+Model Algorithm Less Heterogeneous More Heterogeneous
 REDS[40] Vid4[31] REDS[40] Vid4[31]
-PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
-VRT[25] 16, 35.6 FedAvg[31] 29.71 0.8522 24.88 0.7680 28.85 0.8282 24.60 0.7516
-FedProx[23] 28.82 0.8281 24.18 0.7341 28.64 0.8236 23.99 0.7256
-SCAFFOLD[18] 28.59 0.8181 23.99 0.7182 28.17 0.8074 24.04 0.7208
-FedVSR (Ours) 29.86 0.8562 24.92 0.7682 29.03 0.8339 24.73 0.7584
-RVRT[24] 30, 10.8 FedAvg[31] 31.30 0.8877 25.96 0.8156 29.89 0.8578 25.67 0.8040
-FedProx[23] 30.09 0.8618 25.20 0.7775 29.59 0.8526 25.02 0.7730
-SCAFFOLD[18] 29.63 0.7967 24.79 0.7317 26.71 0.7142 23.08 0.6441
-FedVSR (Ours) 31.49 0.8908 26.05 0.8191 30.34 0.8678 25.90 0.8048
-IART[49] 16, 13.4 FedAvg[31] 30.56 0.8729 25.43 0.7949 29.52 0.8493 25.04 0.7774
-FedProx[23] 29.63 0.8500 24.72 0.7599 29.30 0.8444 24.42 0.7505
-SCAFFOLD[18] 30.29 0.8678 25.18 0.7857 29.26 0.8413 24.85 0.7690
-FedVSR (Ours) 30.88 0.8812 25.54 0.8012 29.70 0.8518 25.26 0.7850
-Table 1: PSNR and SSIM comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+VRT [30] FedAvg[36] 29.71 0.8522 0.1916 24.88 0.7680 0.2446 28.85 0.8282 0.2200 24.60 0.7516 0.2557
+FedProx[27] 28.82 0.8281 0.2133 24.18 0.7341 0.2668 28.64 0.8236 0.2222 24.10 0.7256 0.2713
+SCAFFOLD[21] 28.59 0.8181 0.2228 23.99 0.7182 0.2737 28.17 0.8074 0.2318 24.04 0.7208 0.2739
+FedVSR (Ours) 29.86 0.8562 0.1908 24.92 0.7682 0.2393 29.03 0.8339 0.2123 24.73 0.7584 0.2492
+RVRT [29] FedAvg[36] 31.30 0.8877 0.1536 25.96 0.8156 0.1974 29.89 0.8578 0.1867 25.67 0.8040 0.2065
+FedProx[27] 30.09 0.8618 0.1899 25.20 0.7775 0.2330 29.59 0.8526 0.1943 25.02 0.7730 0.2353
+SCAFFOLD[21] 29.63 0.7967 0.2101 24.79 0.7317 0.2407 26.71 0.7142 0.2590 23.08 0.6441 0.3018
+FedVSR (Ours) 31.54 0.8922 0.1455 26.05 0.8191 0.1889 30.90 0.8811 0.1576 25.90 0.8131 0.1910
+IART [54] FedAvg[36] 30.56 0.8729 0.1695 25.43 0.7949 0.2118 29.52 0.8493 0.1909 25.04 0.7774 0.2272
+FedProx[27] 29.65 0.8500 0.1683 24.72 0.7599 0.2307 29.30 0.8444 0.1958 24.42 0.7505 0.2293
+SCAFFOLD[21] 30.29 0.8678 0.1692 25.18 0.7857 0.2084 29.26 0.8413 0.1992 24.85 0.7690 0.2278
+FedVSR (Ours) 30.88 0.8812 0.1556 25.54 0.8012 0.1990 29.76 0.8536 0.1887 25.26 0.7850 0.2220
+Table 1: PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$) comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
 5.1 setup
@@
 For our baseline, we use a combination of FedAvg[36], FedProx[27], and SCAFFOLD[21] from the FL side, along with VRT[30], RVRT[29], and IART[54] from VSR.
+We selected these FL methods as they are widely accepted and frequently used to benchmark core algorithmic performance in recent literature.
+While newer approaches like MOON[24], FedDyn[1], FedDisco[57], FedLAW[28] exist, they are rely on task-specific assumptions such as logits, class-wise statistics, or server-side data proxies that do not translate well to low-level, regression-based tasks like VSR.
+On the other hand, exisitng FL-SR models such as FedSR, p²FedSR, and FedMRI are not directly comparable to FedVSR for two key reasons.
+First, they are tightly coupled to a specific model architecture: they require dividing the model into predefined blocks (“shallow–deep–upsampler”), which limits flexibility and demands structural changes to existing SR models.
+Second, these methods are stateful meaning each client must store additional data across training rounds, such as degradation descriptors, deep feature extractors, or previous model weights.
+This leads to significantly higher storage, computation, and coordination overhead.
+In contrast, FedVSR is designed to be both model-agnostic and stateless, making it simpler, more lightweight, and easier to apply in diverse real-world scenarios.
 Evaluation Metrics:
-The quality of the reconstructed video is typically evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM).
+The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS).
 Figure 2: PSNR values across different rounds for various test sets under different settings for VRT [30], RVRT [29], and IART [54].
@@
 As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets.
-On average, FedVSR surpasses existing FL methods by a margin of 0.89 dB in PSNR and 0.028 in SSIM in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.80 dB in PSNR and 0.027 in SSIM, demonstrating its robustness under challenging conditions.
+On average, FedVSR surpasses existing FL methods by a margin of 0.61 dB in PSNR and 0.022 in SSIM and 0.022 in LPIPS in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.66 dB in PSNR and 0.028 in SSIM and 0.019 in LPIPS, demonstrating its robustness under challenging conditions.
 Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin.
 In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods.
-When comparing FedVSR only against the second-best FL algorithm, FedAvg, our approach achieves an average PSNR margin of 0.22 dB in less heterogeneous settings and 0.27 dB in more heterogeneous settings.
-Similarly, the SSIM improvement over FedAvg is 0.005 in less heterogeneous cases and 0.006 in more heterogeneous cases.
-The smallest improvement relative to FedAvg is observed for VRT in PSNR and IART in SSIM, while the highest improvement is consistently achieved by RVRT, particularly in more heterogeneous settings, further validating FedVSR’s ability to handle real-world federated VSR challenges effectively.
-The findings emphasize the importance of high-frequency-aware optimization in VSR FL-based training.
 The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results.
@@
 The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases.
-However, we notice that for any $\mu$ greater than $31e-$ 3, the VSR model struggles significantly with learning.
+However, we notice that for any $\mu$ greater than 0.001, the VSR model struggles significantly with learning.
 This suggests that the relationship is more complex and requires further investigation and analysis.
@@
 In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches.
-Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 7.
+Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 8.
 5.3 Ablation Study
@@
 These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights.
+Due to space limitations, LPIPS metric results are provided in the supplementary material 7
 Model Setting Less Heterogeneous More Heterogeneous
@@
 Supplementary Material
+7 Extended Evaluation of HiFr Loss and Adaptive Aggregation Using LPIPS
+To complement the analysis based on PSNR and SSIM discussed earlier in the paper, we present here an additional evaluation using LPIPS, which serves as a perceptual similarity metric.
+This acts as an extension to the previously discussed metrics and provides a more complete picture of model performance.
+Due to space limitations, we include this analysis here.
+To evaluate the impact of the DWT loss term and the loss-aware adaptive aggregation strategies, we conducted a set of ablation experiments using FedAvg as the baseline.
+Each component was added independently to isolate its individual effect.
+Specifically, we assessed the contribution of the $\mathcal{L}_{\text{HiFr}}$ term by applying it to FedAvg (F-$\mathcal{L}_{\text{HiFr}}$), and examined two aggregation variants: a greedy loss-aware approach (F-GLA) that prioritizes updates from clients with lower losses, and an adaptive version (F-ALA) that incorporates a decay factor to regulate client influence over time.
+As shown in Table 3 with LPIPS results, the observed trends remain largely consistent across metrics.
+The DWT loss term consistently provides more significant gains compared to the aggregation strategies, and both methods individually outperform the standard FedAvg baseline.
+However, similar to PSNR and SSIM, the greedy strategy sometimes leads to only marginal improvements or even minor performance degradation, especially under high heterogeneity.
+These findings underscore the importance of high-frequency information not only for boosting perceptual quality but also for reliably identifying valuable client updates.
+Model Setting Less Heterogeneous More Heterogeneous
+REDS Vid4 REDS Vid4
+PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
+VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 0.1910 24.91 0.7681 0.2399 28.98 0.8322 0.2130 24.70 0.7577 0.2500
+F-ALA 29.75 0.8531 0.1912 24.89 0.7681 0.2411 28.90 0.8313 0.2140 24.64 0.7559 0.2520
+F-GLA 29.70 0.8523 0.1914 24.86 0.7670 0.2436 28.83 0.8279 0.2198 24.55 0.7500 0.2555
+RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 0.1506 26.02 0.8178 0.1910 30.24 0.8617 0.1612 25.84 0.8042 0.1978
+F-ALA 31.38 0.8880 0.1518 26.00 0.8171 0.1931 30.19 0.8601 0.1728 25.79 0.8041 0.1992
+F-GLA 31.31 0.8879 0.1527 25.95 0.8155 0.1948 29.88 0.8576 0.1807 25.60 0.8032 0.2011
+IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 0.1588 25.46 0.8001 0.1999 29.64 0.8508 0.1899 25.21 0.7837 0.2231
+F-ALA 30.70 0.8789 0.1591 25.44 0.7998 0.2101 29.60 0.8500 0.1903 25.18 0.7831 0.2246
+F-GLA 30.57 0.8730 0.1690 25.42 0.7951 0.2119 29.50 0.8482 0.1929 24.99 0.7761 0.2284
+Table 3: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR, SSIM, and predicted LPIPS under different heterogeneity settings.
 8 Further Analysis: Performance Metrics and Visual Comparisons
 This section provides additional performance evaluations for our proposed method.
-Figures 3, 4, and 5 illustrate the PSNR and SSIM trends across different rounds for multiple test sets under various configurations of VRT [25], RVRT [24], and IART [49].
+Figures 3, 5, and 7 illustrate the PSNR, SSIM and LPIPS trends across different rounds for multiple test sets under various configurations of VRT [30], RVRT [29], and IART [54].
 These results further highlight the effectiveness of our approach in enhancing the performance across different models.
-Figures 6, 7, and 8 provide qualitative visual comparisons on various test frames.
+Figures 4, 6, and 8 provide qualitative visual comparisons on various test frames.
 Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR).
 The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
-Figure 3: PSNR and SSIM values across different rounds for various test sets under different settings for VRT[25].
-Figure 4: PSNR and SSIM values across different rounds for various test sets under different settings for RVRT[24].
-Figure 5: PSNR and SSIM values across different rounds for various test sets under different settings for IART[49].
-Figure 6: Comparison of RVRT output trained with different Federated Learning algorithms under more heterogeneous setting
-Figure 7: Comparison of VRT output trained with different Federated Learning algorithms under more heterogeneous setting
-Figure 8: Comparison of IART output trained with different Federated Learning algorithms under less heterogeneous setting
+Figure 3: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for VRT.
+Refer to caption (a) Frame 000, Clip 015, REDS
+Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 32.41, 0.8981, 0.1456 Refer to caption (d) FedProx[27] 32.10, 0.8935, 0.1511
+Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 28.85, 0.7625, 0.1934 Refer to caption (g) FedVSR (Ours) 33.79, 0.9169, 0.1151
+Figure 4: Comparison of RVRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
+Figure 5: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for RVRT.
+Refer to caption (a) Frame 040, Clip 1, Vid4
+Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 21.15, 0.7185, 0.3520 Refer to caption (d) FedProx[27] 20.43, 0.6735, 0.4101
+Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 20.41, 0.6679, 0.3803 Refer to caption (g) FedVSR (Ours) 21.19, 0.7200, 0.3433
+Figure 6: Comparison of VRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
+Figure 7: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for IART.
+Refer to caption (a) Frame 070, Clip 000, REDS
+Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 27.87, 0.8211, 0.2228 Refer to caption (d) FedProx[27] 27.54, 0.8132, 0.2247
+Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 27.69, 0.8151, 0.2207 Refer to caption (g) FedVSR (Ours) 28.07, 0.8319, 0.2014
+Figure 8: Comparison of IART output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.

### v3

@@
 Abstract
-Video super-resolution aims to enhance low-resolution videos by leveraging both spatial and temporal information.
+Video super-resolution (VSR) aims to enhance low-resolution videos by leveraging both spatial and temporal information.
 While deep learning has led to impressive progress, it typically requires centralized data, which raises privacy concerns.
-Federated learning offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs.
+Federated learning (FL) offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs.
 To address this, we introduce FedVSR, the first FL framework specifically designed for VSR.
-It is model-agnostic and stateless, and introduces a lightweight loss function based on the DWT to better preserve high-frequency details during local training.
+It is model-agnostic and stateless, and introduces a lightweight loss function based on the Discrete Wavelet Transform (DWT) to better preserve high-frequency details during local training.
 Additionally, a loss-aware aggregation strategy combines both DWT-based and task-specific losses to guide global updates effectively.
-Extensive experiments across multiple VSR models and datasets demonstrate that FedVSR consistently outperforms existing FL methods, achieving up to 0.82 dB higher PSNR, 0.0327 higher SSIM, and 0.0251 lower LPIPS.
-These results underscore FedVSR’s ability to bridge the gap between privacy and performance, setting a new benchmark for federated learning in low-level vision tasks.
+Extensive experiments across multiple VSR models and datasets show that FedVSR not only improves perceptual video quality (up to +0.89 dB PSNR, +0.0370 SSIM, 0.0347-0.0347 LPIPS and 4.98 VMAF) but also achieves these gains with close to zero computation and communication overhead compared to its rivals.
+These results demonstrate FedVSR’s potential to bridge the gap between privacy, efficiency, and perceptual quality, setting a new benchmark for federated learning in low-level vision tasks.
 The code is available at: https://github.com/alimd94/FedVSR
@@
 Refer to caption (a) Frame 029, Clip 011, REDS[35]
-Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 30.41, 0.8600, 0.1777 Refer to caption (d) FedProx[27] 30.07, 0.8531, 0.1851
+Refer to caption (b) FedAvg[30] 30.41, 0.8600, 0.1777 Refer to caption (c) FedMedian [52] 28.89, 0.8224, 0.2106 Refer to caption (d) FedProx[21] 30.07, 0.8531, 0.1851
 Refer to caption (e) GT Refer to caption (f) SCAFFOLD[18] 26.78, 0.7148, 0.2516 Refer to caption (g) FedVSR (Ours) 32.00, 0.8889, 0.1508
@@
 Federated learning (FL) provides a principled remedy to these concerns [30].
-In many VSR scenarios, such as smartphones, surveillance, health-related domains, and remote edge sensors, videos are captured on devices with limited bandwidth and strict privacy constraints.
-Uploading hours of footage can consume tens of gigabytes of storage; for instance, even a short 10-second 8K video at 30 fps in raw format can exceed 10 GB in size [2].
-Additionally, uploading such data may violate policies and laws such as GDPR or HIPAA [19].
+In numerous Video Super-Resolution (VSR) scenarios, such as remote patient monitoring, personal phone footage analysis, and confidential surveillance, video data captured on edge devices is inherently privacy-sensitive.
+Training on this data centrally is often prohibited by stringent regulations such as GDPR or CCPA [25], which forbid uploading and aggregating raw footage containing personally identifiable information.
+Privacy is the core motivator for FL in VSR.
+This decentralization also solves the massive logistical challenge of video data size; for instance, attempting to upload even a short 10-second 8K video for training can easily exceed 10 GB [2].
 Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one.
@@
 These combined benefits make FL a logical and effective foundation for VSR systems that face such concerns.
-Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [21, 27, 51], facial recognition [41, 34], and segmentation [37, 34].
+Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [18, 21, 46], facial recognition [29], and segmentation [31, 29].
 However, their application in super-resolution (SR) remains largely unexplored.
-Existing studies [39, 56] found that directly applying existing general FL techniques to SR tasks results in poor-quality reconstructions with blurred textures, mainly due to their inability to effectively capture rich low-level information such as texture, edges and motion.
-Motivated by these limitations, we propose FedVSR, a novel model-agnostic framework for VSR under FL.
-FedVSR treats VSR models as black boxes, allowing it to work with a variety of architectures without requiring architectural changes or access to intermediate layers.
-To improve performance under FL constraints, we introduce a lightweight loss term based on the Discrete Wavelet Transform (DWT), which emphasizes high-frequency details that are often lost during decentralized training.
-While high-frequency preservation is valuable for all SR tasks, its importance is amplified in VSR due to the added complexity of maintaining temporal coherence across frames.
-In federated settings, data heterogeneity and limited communication intensify this challenge.
-FedVSR addresses these issues without maintaining state between rounds, making it both scalable and practical for privacy-sensitive, resource-constrained environments.
+Existing studies [34, 50] found that directly applying general FL techniques to low-level vision tasks like SR results in poor-quality reconstructions with blurred textures.
+This failure is compounded in VSR, where data heterogeneity (diverse video content and quality) and communication constraints severely degrade the model’s ability to capture and synchronize rich spatial and temporal high-frequency details across frames, a prerequisite for high-fidelity video output.
+Motivated by these limitations, we propose FedVSR, a novel framework for VSR under FL.
+FedVSR is model-agnostic, meaning it treats underlying VSR models as black boxes and works by modifying only the local training loss and the server aggregation strategy.
+This enables compatibility with diverse VSR architectures without requiring internal model modifications or access to intermediate layers, which is crucial for scalability and flexibility in a real-world federated deployment.
+The framework consists of two key components.
+The first component is a novel, frequency-aware training phase enhanced by a lightweight 3D Discrete Wavelet Transform (DWT)-based loss.
+By operating in the frequency domain, this loss forces clients to prioritize the spatial and temporal high-frequency details essential for VSR that are typically lost due to FL’s limited communication budget.
+Crucially, we demonstrate that this loss is specific to the FL environment, as its direct application in centralized VSR leads to performance degradation (see Section 5.4.1).
+The second is an innovative adaptive loss aggregation mechanism that explicitly uses client training quality to mitigate the detrimental effects of data heterogeneity and noisy clients.
+While high-frequency preservation is important for all SR tasks, it is especially crucial in VSR to maintain temporal coherence across frames.
+Federated settings further amplify this challenge due to data heterogeneity and limited communication.
+By combining its two components, FedVSR enhances training and aggregation robustness while remaining scalable, state-free, and practical for privacy-sensitive, resource-constrained environments.
 In summary, the main contributions of this paper are as follows:
-  • We design a frequency-aware loss using DWT to sharpen fine details during client updates, and introduce a complementary aggregation method to weight model contributions more effectively.
-To the best of our knowledge, this is the first algorithm specifically designed to address VSR in the FL setting, effectively bridging the gap between privacy preservation and reconstruction quality.
+  • We design a two-pronged solution to the Federated VSR challenge: (1) a novel, frequency-aware loss using 3D DWT to preserve critical spatial-temporal high-frequency details during decentralized training, and (2) an adaptive loss-aware aggregation method that dynamically adjusts client weights to combat data heterogeneity.
+This combination represents the first holistic framework designed to bridge the gap between VSR quality and FL privacy.
   • Our approach makes no architecture-specific assumptions and does not require storing past weights, making it stateless and model-agnostic.
 This design reduces memory usage, avoids privacy risks, and eliminates synchronization overhead in federated learning on edge devices.
-  • Extensive experiments validate FedVSR’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
+  • Extensive experiments validate FedVSR’s effectiveness and robustness among existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
+Figure 2: Overview of the proposed FedVSR framework.
+Each client computes a model-agnostic VSR update augmented with a DWT-based high-frequency loss.
+Clients also track the average local loss , which is used for loss-aware weighted aggregation at the server.
+The global model is iteratively refined while maintaining model-agnostic and stateless properties.
 2 Related Works
 In this section, we begin by discussing the most significant contributions in general FL and VSR. We then explore studies that specifically focus on applying FL to the SR task.
-2.1 Federated Learning
-Federated learning (FL) is a decentralized learning paradigm that enables multiple clients to collaboratively train models while preserving data privacy.
-It has been widely applied across diverse domains, including healthcare, finance, and IoT, where data security and regulatory constraints are critical [3, 7, 62].
-A foundational approach in this domain is FedAvg [36], which follows a four-step iterative process: (1) the server distributes a global model to all participating clients, (2) each client trains its local model using its own data, (3) the locally updated models are sent back to the central server, and (4) the server aggregates these updates to refine the global model for the next training round.
-To enhance the performance of federated learning, multiple methods focus on optimizing local updates.
-FedProx [27] incorporates a proximal term in the local objective to ensure consistency between local and global models.
-SCAFFOLD [21] introduces control variates for both clients and the server.
-FedNova [51] employs normalized averaging to address objective inconsistencies.
-More recently, MOON [25] applies contrastive learning to assess similarity among different model representations.
-2.2 Video Super Resolution
-Early VSR methods that applied SISR techniques frame-by-frame neglected the temporal context, resulting in temporal inconsistencies, loss of detail, and undesired smoothing in the high-resolution output [33].
+2.1 Video Super-Resolution
+Early VSR methods that applied SISR techniques frame-by-frame neglected the temporal context, resulting in temporal inconsistencies, loss of detail, and undesired smoothing in the high resolution output [28].
 VSR leverages multiple temporally correlated low-resolution frames within a video sequence to generate high-resolution output.
@@
 DiQP [11], the first model for 8K videos, employs Denoising Diffusion to counteract compression artifacts while maintaining robustness across varying levels.
-2.3 Federated Learning on Super Resolution Task
-For low-level vision tasks, FedMRI [15] introduces a federated learning framework for MRI reconstruction with a shared encoder for generalized features and client-specific decoders for personalization.
-FedPR [16] improves image quality using pre-trained models and prompt learning with minimal trainable parameters.
-p²FedSR [55] enhances privacy-preserving SR by integrating client-specific degradation priors and a selective model aggregation strategy for personalized FL.
-FedSR [56] employs a detail-assisted contrastive loss for feature alignment at the client level and a hierarchical aggregation policy at the server level with historical global model, optimizing layer-wise model integration.
-3 Preliminaries
-3.1 VSR Problem Definition
-VSR enhances video resolution by using multiple low-resolution frames, leveraging temporal information to exploit inter-frame dependencies, unlike SISR.
-Problem Formulation
-In addition to the widely used RGB color space, other representations such as the YUV and YCbCr color spaces are frequently employed in VSR applications.
-Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in a given LR video sequence, while $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ represents its corresponding high-resolution (HR) counterpart, where $s$ is the scaling factor (e.g., $s=2,4,8$).
-Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius.
-The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as:
-$I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha})$ (1)
-where $\phi(\cdot;\cdot)$ is the degradation function parameterized by $\theta_{\alpha}$, incorporating various degradation factors such as noise, motion blur, and down-sampling effects.
-A more detailed expression for the degradation process can be written as:
-$I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j}$ (2)
-where $D$ represents the down-sampling operation, $B$ denotes the blurring effect, $E_{i\to j}$ is a warping function that accounts for motion between frames $\hat{I}_{i}$ and $\hat{I}_{j}$, and $n_{j}$ represents the image noise.
-The goal of a VSR algorithm is to reconstruct an HR video sequence from degraded LR frames in a manner that closely resembles the ground truth video.
-This makes this task an ill-posed problem because there exist infinitely many possible HR images $\hat{I}_{i}$ that can produce the same LR image $I_{i}$ after degrading, meaning the degradation operator is highly non-injective [49].
-While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence.
-The inverse process of degradation, which represents the super-resolution task, can be expressed as:
-$\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta})$ (3)
-where $\tilde{I}_{i}$ is the estimated HR frame corresponding to the ground truth $\hat{I}_{i}$, and $\theta_{\beta}$ represents the model parameters involved in the reconstruction process [33].
-3.2 FL Problem Definition
-FL enables multiple clients to train a shared model collaboratively without sharing raw data, enhancing privacy and reducing communication overhead through local computations and periodic model aggregation.
-Problem Formulation
-FL aims to minimize the following objective function across participating clients:
-$\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w})$ (4)
-where:
-  • $K$ is the total number of clients,
-  • $p_{k}$ is the relative weight of client $k$ (typically based on its dataset size),
-  • $F_{k}(\mathbf{w})$ is the local loss function for client $k$,
-  • $\mathbf{w}$ represents the global model parameters.
-Each client updates the global model locally:
-$\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t})$ (5)
-where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme:
-$\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}$ (6)
-Not all clients participate equally in every training round due to data heterogeneity, resource constraints, network instability, or other limitations [59].
-Client averaging based on their contributions stabilizes training, reduces variance, and improves generalization [6].
-However, weighting clients solely by dataset size, as in traditional FedAvg[36], does not fully capture their impact on model performance [44].
-A client’s importance extends beyond data quantity and is also influenced by the quality of its model updates and the loss it generates during training.
-Based on [46], we implement an approach where clients with lower loss are typically better trained and can be assigned higher weights during averaging.
-This ensures improved overall performance by leveraging the robustness properties analyzed in the paper.
-It is worth noting that some works instead prioritize clients with higher loss to increase fairness [26].
-This approach, which contrasts with a focus on final performance, is more common in the context of personalized federated learning.
-By incorporating loss-based importance into the aggregation process, federated learning can achieve more reliable and efficient training.
-This adaptive weighting ensures that client updates more accurately reflect their significance in improving the global model, leading to a more stable and equitable learning process [17].
-4 Methodology
+2.2 Federated Learning
+Federated learning (FL) is a decentralized paradigm that enables multiple clients to collaboratively train models while preserving data privacy.
+It has been widely applied in domains such as healthcare, finance, and the Internet of Things (IoT), where data security and regulatory compliance are critical [3, 7, 57].
+A foundational approach is FedAvg [30], which follows a four-step iterative process: (1) the server distributes a global model to all clients, (2) each client trains its local model on its own data, (3) locally updated models are sent back to the server, and (4) the server aggregates these updates to refine the global model for the next round.
+To improve FL performance, several methods focus on optimizing local updates.
+FedProx [21] adds a proximal term to maintain consistency between local and global models.
+SCAFFOLD [18] introduces control variates for both clients and server, while FedNova [46] employs normalized averaging to address objective inconsistencies.
+More recently, MOON [20] leverages contrastive learning to assess similarity among different model representations.
+Although these methods enhance convergence and stability in classification tasks, their reliance on class-specific assumptions limits applicability to low-level, regression-based tasks such as VSR.
+For instance, MOON, FedDyn [1], FedDisco [51], and FedLAW [22] depend on logits, class-wise statistics, or server-side data proxies—assumptions that do not translate to VSR or SISR scenarios.
+2.3 Federated Learning in Image Super Resolution
+Based on our review of the literature, FL has not yet been explored for VSR; the works discussed below are currently the only FL-based approaches applied to SISR.
+For Image-based low-level vision tasks, FedMRI [14] introduces a federated learning framework for MRI reconstruction using a shared encoder for generalized features and client-specific decoders for personalization.
+FedPR [15] leverages pre-trained models and prompt learning with minimal trainable parameters to improve image quality.
+p²FedSR [49] enhances personalized privacy-preserving SR through client-specific degradation priors and selective model aggregation, while FedSR [50] employs detail-assisted contrastive loss for feature alignment and a hierarchical aggregation policy for layer-wise model integration.
+Although these approaches represent significant progress in FL for low-level vision, they are restricted to SISR rather than VSR.
+Moreover, methods such as FedSR, p²FedSR, and FedMRI are closely tied to specific model architectures (shallow——deep——upsampler), limiting applicability to a narrow range of SR frameworks.
+They also rely on stateful training: clients must retain auxiliary information such as degradation priors, feature statistics, or historical checkpoints, which introduces non-trivial storage, computation, and synchronization overhead.
+Such constraints hinder scalability and deployment, especially in heterogeneous or resource-limited environments.
+In this paper, we propose FedVSR, a model-agnostic and stateless framework that allows any VSR model to be seamlessly incorporated without structural modifications or client-side state management.
+This design ensures easy extensibility, offering greater flexibility and scalability than prior work, and making FedVSR practical for real-world deployment where lightweight, generalizable solutions are essential.
+3 Overview of FedVSR
+FedVSR integrates Video Super-Resolution (VSR) with Federated Learning (FL), enabling collaborative enhancement of video resolution across distributed clients while preserving data privacy.
+This section provides a concise overview of the underlying VSR and FL problems and their formulations.
+3.1 Video Super-Resolution (VSR)
+VSR aims to enhance video resolution by leveraging temporal information across multiple low-resolution (LR) frames, exploiting inter-frame dependencies that single-image super-resolution (SISR) methods cannot utilize:
+Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in an LR video sequence, and $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ its high-resolution (HR) counterpart, where $s$ is the upscaling factor.
+Define a temporal window of $2N+1$ HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$.
+The LR frames are generated via a degradation process:
+$I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha}),$ (1)
+where $\phi(\cdot;\cdot)$ models degradation factors such as blur, noise, and motion. A more explicit formulation is:
+$I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j},$ (2)
+with $D$ as the down-sampling operator, $B$ the blur effect, $E_{i\to j}$ a warping function for motion compensation, and $n_{j}$ image noise. VSR seeks to reconstruct HR frames:
+$\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta}),$ (3)
+where $\tilde{I}_{i}$ approximates the ground truth $\hat{I}_{i}$ [44, 28]. This is an ill-posed problem due to the non-injective nature of the degradation function.
+3.2 Federated Learning (FL)
+FL allows multiple clients to collaboratively train a shared model without exchanging raw data, reducing privacy risks and communication overhead. The global optimization problem is:
+$\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w}),$ (4)
+where $K$ is the number of clients, $p_{k}$ is the relative weight of client $k$, $F_{k}(\mathbf{w})$ the local loss, and $\mathbf{w}$ the global model parameters. Clients update locally via:
+$\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t}),$ (5)
+and the server aggregates updates with weighted averaging:
+$\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}.$ (6)
+Client participation can vary due to data heterogeneity, resource limits, or network instability, and simple dataset-size weighting may not capture each client’s true impact [54, 5, 30, 38].
+FedVSR incorporates hybrid adaptive loss-based weighting, where clients with lower local loss contribute more effectively to the global model, improving training stability and performance.
+See Section 4.2
+3.3 FedVSR: Integrating VSR with FL
+FedVSR combines these paradigms by distributing the VSR training process across multiple clients.
+Each client improves HR frame reconstruction locally while preserving its private video data.
+The server aggregates model updates using adaptive weights to optimize global performance.
+This approach maintains temporal coherence in videos, leverages multi-client data diversity, and preserve privacy to address the challenges of centralized VSR training.
+4 Proposed FedVSR Framework
 VSR remains a challenging problem due to its ill-posed nature.
@@
 Gradient mismatches [47] and parameter inconsistencies slow convergence [10].
-Most FL algorithms are designed for image classification which can be considered as a well-posed problem [18].
-However, VSR is an ill-posed task, requiring a further fundamental assessment of existing methods to ensure effective adaptation.
-For instance, FedProx assumes a certain degree of similarity across client data distributions, but in real-world applications, these distributions can vary significantly.
-When this assumption fails, models experience drift, leading to inconsistent local updates and poor generalization [13].
-Similarly, SCAFFOLD assumes unbiased client updates, yet in large-scale deployments, heterogeneous data causes significant divergence, slowing convergence.
-Moreover, its need for additional state information per client makes it impractical for large-scale FL systems due to communication and storage constraints [10].
-Meanwhile, Other methods such as FedDisco[58] and MOON rely on class label distributions or logits, making them inherently task-specific rather than generalizable.This work proposes a solution to enhance the overall objective of VSR models while improving aggregation to mitigate these challenges.
-4.1 Enhancing Loss Function for Federated VSR
-Selecting a proper loss function as objective for VSR model is crucial and determine the overall output’s performance [50, 47] and as depicted in Fig.1.
+Our design is a two-pronged approach that enhances client training via a novel loss term and improves server robustness through adaptive aggregation.
+The overall architecture of FedVSR framework is illustrated in Fig.
+2
+4.1 High-Frequency Loss for Federated VSR
+Selecting a proper loss function as objective for VSR model is crucial and determine the overall output’s performance [45, 41] and as depicted in Fig.
+1.
 It is evident that existing FL algorithms has problems in capturing and generating details in the output.
 Since we aim to avoid adding computational overhead to the client by using adversarial or perceptual loss, and because our objective is to keep the model untouched, ensuring FedVSR remains model-agnostic, we introduce an additional term to the overall loss of model improve performance.
-4.1.1 DWT for High-Frequency Preservation
+4.1.1 3D DWT for High-Frequency Preservation
 One common failure in training VSR models is producing overly smooth outputs, meaning the trained model fails to capture high-frequency information.
@@
 Since the models we used are among the top-performing ones in a centralized setting, this suggests that the failure to restore high-frequency details is specifically related to training under the FL paradigm.
-To alleviate this, we leverage the Discrete Wavelet Transform (DWT), which has proven effective in super-resolution tasks [20, 61].
+To alleviate this, we leverage the 3D Discrete Wavelet Transform (DWT).
+While 2D DWT has shown effectiveness in SISR [17, 56], we apply the 3D variant to capture temporal high-frequency details across frames in addition to spatial ones.
+This is essential for maintaining temporal coherence in VSR output, a problem standard FL aggregation fails to address effectively.
 DWT is a multi-resolution signal decomposition technique that represents a signal in terms of approximation and detail coefficients at different scales.
-For a 2D signal (such as an image or video frame), DWT decomposes the signal into sub-bands of different frequencies and orientations such as:
-  • Approximation Coefficients $F_{i}^{LL}$ (low-pass filtered in both dimensions).
-  • Horizontal Detail Coefficients $F_{i}^{LH}$ (high-pass in vertical, low-pass in horizontal).
-  • Vertical Detail Coefficients $F_{i}^{HL}$ (low-pass in vertical, high-pass in horizontal).
-  • Diagonal Detail Coefficients $F_{i}^{HH}$ (high-pass in both dimensions).
-One of the simplest wavelets used in DWT is Haar wavelet, which consists of a step function.
-It is computationally efficient and commonly used in image and video processing tasks.
-In this method each frame’s pixel values are convolved with:
-  • Low-pass filter $h=\frac{1}{\sqrt{2}}[1,1]$ to extract approximation coefficients.
-  • High-pass filter $g=\frac{1}{\sqrt{2}}[1,-1]$ to extract detail coefficients.
-This process is repeated independently for each frame in the sequence.
-The coefficient in those four sub-bands are calculated as:
-$F_{i}^{LL}=(F_{i}*h^{T})*h,\quad F_{i}^{LH}=(F_{i}*h^{T})*g$ (7)
-$F_{i}^{HL}=(F_{i}*g^{T})*h,\quad F_{i}^{HH}=(F_{i}*g^{T})*g$ (8)
-The low-frequency component $F_{i}^{LL}$, contains the content and color information of the input frame, while the three high-frequency sub-bands capture detailed information about global structures and textures [43].
-Therefore, for our loss helper term, we consider only the high-frequency sub-bands and ignore the low-frequency ones.
-The loss is then computed pixel-wise in each sub-band between the predicted $\bar{F}$ and ground truth $\hat{F}$ frame at time step $i$:
-$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH}$ (9)
-$L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL}$ (10)
-$L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (11)
-Then with the Charbonnier loss [9] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ :
-$\mathcal{L}_{\text{HiFr}}=\sum_{i}\sum_{s\in\{LH,HL,HH\}}\sqrt{{L_{i}^{s}}^{2}+\epsilon^{2}}$ (12)
-Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows:
+For our implementation, we select the separable Haar wavelet.
+While more sophisticated wavelets offer better decorrelation, Haar’s inherent simplicity and use of short integer filters result in minimal computational overhead during the local update step.
+This is a critical design choice for FL, as it aligns with our goal of introducing a lightweight, non-intensive loss term that does not significantly burden the resource-constrained client devices.
+For a 3D signal (such as a volumetric image or a video clip), DWT decomposes the signal into eight sub-bands of different frequencies and orientations:
+  • Low-Frequency Approximation Coefficients $F_{i}^{LLL}$ (low-pass filtered along all three dimensions).
+  • High-Frequency Coefficients with single high-pass filtering:
+    • $F_{i}^{LLH}$ (low-pass depth and height, high-pass width),
+    • $F_{i}^{LHL}$ (low-pass depth and width, high-pass height),
+    • $F_{i}^{HLL}$ (low-pass height and width, high-pass depth).
+  • High-Frequency Coefficients with two high-pass filterings:
+    • $F_{i}^{LHH}$ (low-pass depth, high-pass height and width),
+    • $F_{i}^{HLH}$ (low-pass height, high-pass depth and width),
+    • $F_{i}^{HHL}$ (low-pass width, high-pass depth and height).
+  • Full High-Frequency Coefficients $F_{i}^{HHH}$ (high-pass in all three dimensions).
+Using the separable Haar wavelet, the 1D filters are defined as:
+  • Low-pass filter $h=\frac{1}{\sqrt{2}}[1,1]$,
+  • High-pass filter $g=\frac{1}{\sqrt{2}}[1,-1]$.
+The coefficients in the eight sub-bands are computed by successive filtering along the three dimensions (depth, height, width). For example:
+$F_{i}^{LLL}$ $=((F_{i}*h^{T})*h)*h$ (7)
+$F_{i}^{LLH}$ $=((F_{i}*h^{T})*h)*g$ (8)
+$F_{i}^{LHL}$ $=((F_{i}*h^{T})*g)*h$ (9)
+$F_{i}^{LHH}$ $=((F_{i}*h^{T})*g)*g$ (10)
+$F_{i}^{HLL}$ $=((F_{i}*g^{T})*h)*h$ (11)
+$F_{i}^{HLH}$ $=((F_{i}*g^{T})*h)*g$ (12)
+$F_{i}^{HHL}$ $=((F_{i}*g^{T})*g)*h$ (13)
+$F_{i}^{HHH}$ $=((F_{i}*g^{T})*g)*g.$ (14)
+Similar to the 2D case, the low-frequency component $F_{i}^{LLL}$ captures the global structure and coarse content of the input, while the seven high-frequency sub-bands capture edges, textures, and structural variations along different orientations [37].
+For our loss design, we focus on the high-frequency sub-bands and ignore the low-frequency approximation. The high-frequency helper loss is then computed using the Charbonnier loss [9]:
+$\mathcal{L}_{\mathrm{HiFr}}(\bar{F},\hat{F})=\mathbb{E}\!\left[\sqrt{\big(H_{\text{HF}}(\bar{F})-H_{\text{HF}}(\hat{F})\big)^{2}+\epsilon^{2}}\right]$ (15)
+where $H_{\text{HF}}(\cdot)$ denotes the concatenation of those seven high frequency sub-bands $\{LLH,LHL,LHH,HLL,HLH,HHL,HHH\}$ obtained by the 3D DWT from the predicted $\bar{F}$ and ground truth $\hat{F}$ frame , $\epsilon>0$ is the Charbonnier constant, and $\mathbb{E}[\cdot]$ indicates the mean over all elements (batch, channels, temporal, and spatial).
+Finally, the total training objective combines the standard VSR loss with the DWT-based high-frequency loss:
 $\mathcal{L}_{\text{total}}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$ (16)
-As our experiments demonstrate (see Section 5.3), the optimal weight values are unbounded, meaning they are not constrained by the relationship $\lambda_{\text{VSR}}=1-\lambda_{\text{HiFr}}$ and can be independently adjusted for better performance.
-4.2 Loss-Aware Aggregation for Federated VSR
-To further enhance the model’s performance, we employ a greedy aggregation approach with a decay rate, inspired by MKL-SGD [46].
-MKL-SGD improves robustness by selecting samples with the lowest loss during training, thereby mitigating the impact of outliers.
-Similarly, in our method, clients accumulate their loss values throughout local training and compute an average loss based on the number of training steps.
-When sending their updated model weights to the server, they also transmit this average loss.
-At the server, we perform inverse weighted averaging, where each client’s weight is computed as $I_{i}=\frac{1}{\mathcal{L}_{i}}$ and adjusted using an adaptive step: $I^{\prime}_{i}=I_{i}^{\alpha}$ where $\alpha$ controls the strength of loss-based weighting.
-To regulate this process, we introduce a decay rate that gradually reduces the effect of loss-based weighting over time, following:$\alpha=\alpha^{1-\frac{t}{T}}$ where $t$ is the current round and $T$ is the total number of rounds.
-This ensures that early training benefits from loss-based prioritization, while later rounds transition smoothly toward standard averaging, balancing robustness and generalization.
+4.2 Adaptive Loss-Aware Aggregation
+To further enhance the model’s performance, we employ a hybrid aggregation strategy that dynamically balances uniform and loss-based weighting.
+In general VSR research, the standard task loss (e.g., L1 or L2) is already widely accepted as a reliable pixel-wise measure of reconstruction quality [28].
+This provides a crucial distinction from federated classification tasks.
+The total local loss ($\mathcal{L}_{\text{total}}$) used in FedVSR maintains this reliability, being calculated as the combination of the standard VSR loss and our novel $\mathcal{L}_{\text{HiFr}}$ high-frequency loss.
+Our $\mathcal{L}_{\text{HiFr}}$ term acts as an enhancement, sensitizing $\mathcal{L}_{\text{total}}$ to the high-frequency details often lost through averaging, thus making the total loss an even better measure of performance.
+Unlike high-level tasks such as classification where a high loss might signal valuable Non-IID data [53] , a high value for $\mathcal{L}_{\text{total}}$ in VSR directly signifies a failure to reconstruct high-fidelity details and temporal coherence, the fundamental objective.
+Therefore, clients with significantly higher $\mathcal{L}_{\text{total}}$ pixel-wise losses genuinely contribute lower-quality model updates, making prioritization justifiable.
+This aggregation strategy, inspired by MKL-SGD [40], improves robustness by prioritizing updates that achieve superior local performance.
+Our method first computes the average local loss ($\mathcal{L}_{\text{total}}$) for each client after training, which is then transmitted to the server as a single scalar value alongside the model weights.
+4.2.1 Deriving Loss-Based Prioritization
+Rather than using inverse loss weights directly, we normalize them and apply an adaptive step parameter to create the loss-based distribution ($l_{i}$), which naturally emphasizes clients with lower losses.
+4.2.2 Hybrid Weighting for Robustness
+A key design feature is that we do not rely on $l_{i}$ alone.
+The aggregation is governed by a convex combination of $l_{i}$ and the uniform distribution ($u_{i}$).
+This $\mathbf{u_{i}}$ component ensures that all clients contribute a baseline minimum update, thereby preventing the complete marginalization of updates derived from challenging, yet strategically important, diverse video data.
+4.2.3 Thresholded Adaptation via Hellinger Distance
+To avoid overreacting to minor variations in client performance, we quantify the divergence between the uniform ($u_{i}$) and loss-based ($l_{i}$) distributions using the Hellinger distance [16].
+This distance serves as a mixing coefficient ($m$).
+By applying a soft threshold ($\tau$) to this distance, the loss-based prioritization is only significantly activated when there is a large, meaningful divergence in client update quality.
+If losses are similar, the aggregation gradually defaults to uniform averaging.
+Finally, the client’s weight ($w_{i}$) in the global update is the convex combination of these two distributions, which is then applied to the model parameters during aggregation.
+This approach allows the aggregation process to adaptively leverage loss-based prioritization when needed, while safeguarding against the unfair penalization of diverse data, thereby improving both robustness and generalization.
 The overall algorithm is summarized in Algorithm 1.
-Algorithm 1 Our Method
-1:Number of communication rounds $T$, Number of clients $N$, Learning rate $\eta$, Adaptive step $\alpha$, Decay rate $\gamma$
-2:Global model weights $W_{T}$
-3:Server Initialization Initialize global model $W_{0}$
+Algorithm 1 Our Method (FedVSR)
+1:Number of communication rounds $T$, Number of clients NN, Learning rate $\eta$, Adaptive step $\alpha$, Decay rate $\gamma$
+2:Global model weights $W^{T}$
+3:Server Initialization: Initialize global model $W^{0}$
 4:for $t=1,\dots,T$ do
@@
 6:  for each client $i\in S_{t}$ in parallel do
-7:        Client updates local model:
-8:         Retrieve latest global model $W_{t}$
-9:         Perform local training for $E$ epochs with learning rate $\eta$
-10:        Compute Local Loss:
-11:         Compute model-specific loss $\mathcal{L}_{\text{VSR}}$
-12:         Compute $\mathcal{L}_{\text{DWT}}$ loss $\mathcal{L}_{\text{DWT}}$
-13:         Compute total loss
-$\mathcal{L}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$
-14:        Client updates local model:
-$\mathbf{w}_{i}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{i}(\mathbf{w}^{t})$
-15:         Return updated weights $W_{i}$ and loss value $\mathcal{L}_{i}$ to server
-16:    end for
-17:    Server Aggregation: (Performed at the Server)
-18:     Compute inverse loss weighting $I_{i}=\frac{1}{\mathcal{L}_{i}}$
-19:     Apply adaptive step $I^{\prime}_{i}=I_{i}^{\alpha}$
-20:     Normalize inverse loss weights $w_{i}=\frac{I^{\prime}_{i}}{\sum_{j\in S_{t}}I^{\prime}_{j}}$
-21:     Compute weighted model update $W_{i}^{\text{scaled}}=w_{i}\cdot W_{i}$
-22:     Aggregate global model $W_{t+1}=\sum_{i\in S_{t}}W_{i}^{\text{scaled}}$
-23:    Apply Decay Rate to Adaptive Step $\alpha=\alpha^{1-\frac{t}{T}}$
-24:end for
-25:return Global model $W_{T}$
+7:    Client downloads latest global model $W^{t}$
+8:    Perform local training for $E$ epochs with learning rate $\eta$
+9:    Compute local losses:
+10:       $\mathcal{L}_{\text{VSR}}$ (task-specific loss)
+11:       $\mathcal{L}_{\text{HiFr}}$ (high-frequency loss)
+12:    Compute total loss:
+$\mathcal{L}_{i}=\lambda_{\text{VSR}}\mathcal{L}_{\text{VSR}}+\lambda_{\text{HiFr}}\mathcal{L}_{\text{HiFr}}$
+13:    Update local model:
+$W_{i}^{t+1}=W^{t}-\eta\nabla F_{i}(W^{t})$
+14:    Return $W_{i}^{t+1}$ and loss value $\mathcal{L}_{i}$ to server
+15:  end for
+16:  Server Aggregation:
+17:  Compute uniform weights $u_{i}=\frac{1}{|S_{t}|}$
+18:  Compute inverse-loss weights $I_{i}=\frac{1}{\mathcal{L}_{i}}$, apply adaptive step $I^{\prime}_{i}=I_{i}^{\alpha}$
+19:  Normalize loss-based weights $l_{i}=\frac{I^{\prime}_{i}}{\sum_{j\in S_{t}}I^{\prime}_{j}}$
+20:  Compute Hellinger distance:
+$H=\sqrt{\tfrac{1}{2}\sum_{i\in S_{t}}(\sqrt{u_{i}}-\sqrt{l_{i}})^{2}}$
+21:  Compute mixing coefficient:
+$m=\begin{cases}0&H<\tau\\ \tfrac{H-\tau}{1-\tau}&H\geq\tau\end{cases}$
+22:  Final aggregation weights: $w_{i}=(1-m)u_{i}+ml_{i}$
+23:  Update global model:
+$W^{t+1}=\sum_{i\in S_{t}}w_{i}\cdot W_{i}^{t+1}$
+24:  Apply Decay Rate to Adaptive Step $\alpha=\alpha^{1-\frac{t}{T}}$
+25:end for
+26:return Global model $W^{T}$
 5 Experiments
-Model Algorithm Less Heterogeneous More Heterogeneous
-REDS[40] Vid4[31] REDS[40] Vid4[31]
-PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-VRT [30] FedAvg[36] 29.71 0.8522 0.1916 24.88 0.7680 0.2446 28.85 0.8282 0.2200 24.60 0.7516 0.2557
-FedProx[27] 28.82 0.8281 0.2133 24.18 0.7341 0.2668 28.64 0.8236 0.2222 24.10 0.7256 0.2713
-SCAFFOLD[21] 28.59 0.8181 0.2228 23.99 0.7182 0.2737 28.17 0.8074 0.2318 24.04 0.7208 0.2739
-FedVSR (Ours) 29.86 0.8562 0.1908 24.92 0.7682 0.2393 29.03 0.8339 0.2123 24.73 0.7584 0.2492
-RVRT [29] FedAvg[36] 31.30 0.8877 0.1536 25.96 0.8156 0.1974 29.89 0.8578 0.1867 25.67 0.8040 0.2065
-FedProx[27] 30.09 0.8618 0.1899 25.20 0.7775 0.2330 29.59 0.8526 0.1943 25.02 0.7730 0.2353
-SCAFFOLD[21] 29.63 0.7967 0.2101 24.79 0.7317 0.2407 26.71 0.7142 0.2590 23.08 0.6441 0.3018
-FedVSR (Ours) 31.54 0.8922 0.1455 26.05 0.8191 0.1889 30.90 0.8811 0.1576 25.90 0.8131 0.1910
-IART [54] FedAvg[36] 30.56 0.8729 0.1695 25.43 0.7949 0.2118 29.52 0.8493 0.1909 25.04 0.7774 0.2272
-FedProx[27] 29.65 0.8500 0.1683 24.72 0.7599 0.2307 29.30 0.8444 0.1958 24.42 0.7505 0.2293
-SCAFFOLD[21] 30.29 0.8678 0.1692 25.18 0.7857 0.2084 29.26 0.8413 0.1992 24.85 0.7690 0.2278
-FedVSR (Ours) 30.88 0.8812 0.1556 25.54 0.8012 0.1990 29.76 0.8536 0.1887 25.26 0.7850 0.2220
-Table 1: PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$) comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
+Model Params (M) Algorithm Less Heterogeneous
+REDS [35] Vid4 [26]
+[…]
+VRT [24] 35.6 FedAvg [30] 29.71 0.8522 0.1916 96.59 24.88 0.7680 0.2446 81.62
+FedProx [21] 28.82 0.8281 0.2133 92.95 24.18 0.7341 0.2668 76.03
+SCAFFOLD [18] 28.59 0.8181 0.2228 91.45 23.99 0.7182 0.2737 73.15
+FedMedian [52] 27.89 0.7999 0.2379 87.46 23.59 0.6997 0.2963 70.62
+FedVSR (Ours) 29.96 0.8595 0.1807 97.47 24.97 0.7745 0.2284 83.07
+RVRT [23] 10.8 FedAvg [30] 31.30 0.8877 0.1536 97.97 25.96 0.8156 0.1974 86.05
+FedProx [21] 30.09 0.8618 0.1899 95.65 25.20 0.7775 0.2330 82.01
+SCAFFOLD [18] 29.63 0.7967 0.2101 94.71 24.79 0.7317 0.2407 81.01
+FedMedian [52] 29.76 0.8444 0.1931 94.79 24.88 0.7602 0.2404 80.11
+FedVSR (Ours) 31.64 0.8956 0.1377 98.94 26.10 0.8207 0.1802 87.37
+IART [48] 13.4 FedAvg [30] 30.56 0.8729 0.1695 97.84 25.43 0.7949 0.2118 84.34
+FedProx [21] 29.65 0.8500 0.1683 96.14 24.72 0.7599 0.2307 80.65
+SCAFFOLD [18] 30.29 0.8678 0.1692 97.47 25.18 0.7857 0.2084 83.38
+FedMedian [52] 29.17 0.8265 0.1960 95.55 24.40 0.7447 0.2302 78.50
+FedVSR (Ours) 30.98 0.8845 0.1473 98.45 25.59 0.8077 0.1899 86.12
+Table 1: Comparison of PSNR ($\uparrow$), SSIM ($\uparrow$), LPIPS ($\downarrow$), and VMAF ($\uparrow$) for different FL algorithms under less heterogeneous data distributions across VSR models and datasets.
+Model Params (M) Algorithm More Heterogeneous
+REDS [35] Vid4 [26]
+[…]
+VRT [24] 35.6 FedAvg [30] 28.85 0.8282 0.2200 91.81 24.60 0.7516 0.2557 77.52
+FedProx [21] 28.64 0.8236 0.2222 88.04 24.10 0.7256 0.2713 72.16
+SCAFFOLD [18] 28.17 0.8074 0.2318 87.83 24.04 0.7208 0.2739 71.85
+FedMedian [52] 27.68 0.7761 0.2429 85.72 23.77 0.7001 0.2890 69.69
+FedVSR (Ours) 29.12 0.8371 0.2010 94.37 24.78 0.7646 0.2378 80.10
+RVRT [23] 10.8 FedAvg [30] 29.89 0.8578 0.1867 96.31 25.67 0.8040 0.2065 84.13
+FedProx [21] 29.59 0.8526 0.1943 94.71 25.02 0.7730 0.2353 82.01
+SCAFFOLD [18] 26.71 0.7142 0.2590 72.38 23.08 0.6441 0.3018 57.54
+FedMedian [52] 28.85 0.8202 0.2119 87.84 24.51 0.7419 0.2499 74.73
+FedVSR (Ours) 31.00 0.8844 0.1492 97.45 25.95 0.8197 0.1823 85.85
+IART [48] 13.4 FedAvg [30] 29.52 0.8493 0.1909 95.99 25.04 0.7774 0.2272 81.71
+FedProx [21] 29.30 0.8444 0.1958 94.16 24.42 0.7505 0.2293 77.42
+SCAFFOLD [18] 29.26 0.8413 0.1992 94.16 24.85 0.7690 0.2278 79.40
+FedMedian [52] 28.80 0.8280 0.2330 88.90 23.88 0.7350 0.2555 73.99
+FedVSR (Ours) 29.85 0.8568 0.1787 96.84 25.31 0.7914 0.2118 83.18
+Table 2: Comparison of PSNR ($\uparrow$), SSIM ($\uparrow$), LPIPS ($\downarrow$), and VMAF ($\uparrow$) for different FL algorithms under more heterogeneous data distributions across VSR models and datasets.
 5.1 setup
-Dataset:
-Following [30, 29, 54], one of the most commonly used datasets for training VSR is REDS [40].
-It contains 240, 30, and 30 sequences for training, validation, and testing, respectively, with each sequence consisting of 100 frames at a resolution of 720 × 1280.
-The corresponding low-resolution (LR) frames are generated through bicubic downsampling from the high-resolution (HR) images.
-This dataset features a variety of natural and handmade objects, locations, and diverse content, making it heterogeneous in decentralized settings [40].
-To further evaluate FedVSR and assess its robustness in more realistic, highly heterogeneous scenarios, we conducted additional experiments using Kinetics-400 [22], a dataset primarily used for human action recognition.
-It includes 400 action classes with 306,245 videos, each approximately 10 seconds long, sourced from unique YouTube videos.
-Since Kinetics-400 contains videos of varying resolutions, we selected only those with a 720 × 1280 resolution to ensure a fair comparison with REDS.
-From the 400 classes, we randomly chose 40, and for each class, selected six videos, extracting 100-frame clips to match REDS’ format.
-Kinetics-400 is particularly well-suited for testing FedVSR under more extreme heterogeneity because, unlike professionally filmed datasets, its videos often exhibit camera shake, lighting variations, shadows, and background clutter [22].
-This makes it a challenging but realistic benchmark for assessing the model’s adaptability in diverse and less controlled environments.
-For consistency with our experimental setup, we refer to REDS as the ”less heterogeneous” dataset and Kinetics-400 as the ”more heterogeneous” dataset, given the increased variability in content, motion, and environmental conditions in the latter.
-Implementation Details:
+5.1.1 Dataset:
+We designed two distinct training environments to comprehensively evaluate the robustness and generalization of FedVSR under varying levels of data heterogeneity, utilizing two foundational video datasets: REDS [35] and Kinetics-400 [6].
+The REDS dataset is high-quality and professionally recorded, designed specifically for video restoration tasks, offering high content diversity but remaining relatively controlled, with sequences at $720\times 1280$ resolution.
+In contrast, Kinetics-400,dataset primarily used for human action recognition, is a massive dataset sourced from unique amateur YouTube videos, whose content is highly realistic and exhibits significant camera shake, lighting variations, and low control, making it ideal for simulating highly diverse, uncontrolled, edge-device data.
+For the Less Heterogeneous Scenario, we utilize the training sequences of REDS, splitting the data randomly ($\mathbf{IID}$-like) among clients to simulate a less diverse video environment.
+For the More Heterogeneous Scenario, we utilize a custom split of Kinetics-400 to simulate a strong $\mathbf{Non\text{-}IID}$ environment.
+Critically, because VSR does not use labels and developing an natural metric for video heterogeneity would be out of scope, we exploit the existing human action labels (40 classes selected) of Kinetics-400 solely to partition the video clips into Non-IID client silos, maximizing the distribution shift between clients.
+Only $720\times 1280$ videos are selected from Kinetics-400 for format consistency with REDS.
+To assess the generalization capability of the models trained under both scenarios, we use two standard external test sets: the REDS Test set and the Vid4 dataset [26].
+The models trained on both the REDS and Kinetics-400 splits are consistently evaluated against both benchmarks.
+The reason for selecting these specific datasets for testing is to strictly adhere to VSR community standards for better comparison of performance and measurement, since they are established benchmarks.
+5.1.2 Implementation Details:
 To maintain our objective of developing a model-agnostic framework, we did not alter the official parameters of the baseline VSR models.
 Instead, we used their recommended settings for training.
-We created 40 clients, each assigned six videos in both experimental setups.
+We created 40 clients, with participation of 10% in both experimental setups.
 In the more heterogeneous setup, each client received videos from a single class; for example, one client had only ”archery” videos, while another had only ”making pizza” videos.
@@
 We set the total number of global rounds to 100, with each client performing one local round (epoch) per global round.
-In each round, we randomly selected four clients to perform training.
 After completing their local training, these clients sent their updated weights to the server for aggregation.
 All experiments were conducted using bi-cubic (BI) degradation.
-Baselines:
-For our baseline, we use a combination of FedAvg[36], FedProx[27], and SCAFFOLD[21] from the FL side, along with VRT[30], RVRT[29], and IART[54] from VSR.
+5.1.3 Baselines:
+For our baseline, we use a combination of FedAvg [30], FedProx [21], FedMedian [52], and SCAFFOLD [18] from the FL side, along with VRT [24], RVRT [23], and IART [48] from VSR.
 We selected these FL methods as they are widely accepted and frequently used to benchmark core algorithmic performance in recent literature.
-While newer approaches like MOON[24], FedDyn[1], FedDisco[57], FedLAW[28] exist, they are rely on task-specific assumptions such as logits, class-wise statistics, or server-side data proxies that do not translate well to low-level, regression-based tasks like VSR.
-On the other hand, exisitng FL-SR models such as FedSR, p²FedSR, and FedMRI are not directly comparable to FedVSR for two key reasons.
-First, they are tightly coupled to a specific model architecture: they require dividing the model into predefined blocks (“shallow–deep–upsampler”), which limits flexibility and demands structural changes to existing SR models.
-Second, these methods are stateful meaning each client must store additional data across training rounds, such as degradation descriptors, deep feature extractors, or previous model weights.
-This leads to significantly higher storage, computation, and coordination overhead.
-In contrast, FedVSR is designed to be both model-agnostic and stateless, making it simpler, more lightweight, and easier to apply in diverse real-world scenarios.
-Evaluation Metrics:
-The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS).
-Figure 2: PSNR values across different rounds for various test sets under different settings for VRT [30], RVRT [29], and IART [54].
-Reproducibility:
+5.1.4 Evaluation Metrics:
+The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS) and Video Multimethod Assessment Fusion (VMAF).
+5.1.5 Reproducibility:
 To ensure our experiments are reproducible, we fixed the random seed, which doesn’t affect client selection, model initialization, etc.
 We used four NVIDIA A6000 GPUs for training, with each model under a federated learning algorithm taking an average 46 hours to train.
-For the baselines alone, this consumed 36 days (2 experiments × 9 models+algorithms × 2 days per run).
+For the baselines alone, this consumed 1102 hours (2 experiments × 3 VSR models × 4 Fl algorithms × 46 hours per run).
 5.2 Results and Comparison
-Quantitative Results:
-As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets.
-On average, FedVSR surpasses existing FL methods by a margin of 0.61 dB in PSNR and 0.022 in SSIM and 0.022 in LPIPS in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.66 dB in PSNR and 0.028 in SSIM and 0.019 in LPIPS, demonstrating its robustness under challenging conditions.
+5.2.1 Quantitative Results:
+As shown in Tables  1 and  2, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets.
+On average, FedVSR surpasses existing FL methods by a margin of 0.82 dB in PSNR and 0.0337 in SSIM and 0.0334 in LPIPS and 3.62 in VMAF in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.96 dB in PSNR and 0.0404 in SSIM and 0.0359 in LPIPS and 6.34 in VMAF, demonstrating its robustness under challenging conditions.
 Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin.
 In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods.
-The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results.
-As shown in Fig.2, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FedAvg, FedProx, and SCAFFOLD, particularly in later training rounds.
+The observed performance trends in our experiments align with existing findings on FedProx, SCAFFOLD and FedMedian further validating our results.
+As shown in Fig.
+3, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FL baselines particularly in later training rounds.
 This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions.
@@
 This suggests that the relationship is more complex and requires further investigation and analysis.
-Similarly, our findings regarding SCAFFOLD’s inefficacy in highly heterogeneous settings corroborate previous research that highlights its susceptibility to instability under heterogeneous data [4].
-The significant performance drops and stagnation observed in our experiments, particularly with the RVRT model, align with the reported issues of SCAFFOLD transmitting nearly twice the data of other algorithms due to its reliance on control variates.
-This added communication burden does not necessarily translate to better convergence, as our results show that SCAFFOLD struggles to optimize effectively in diverse settings.
-Prior work [4] also notes that SCAFFOLD is highly prone to sudden performance breakdowns when gradient clipping is not employed, which is reflected in our results where its PSNR fluctuates or plateaus early in highly heterogeneous scenarios.
+Figure 3: PSNR across different rounds for various test sets under different settings for VRT [24], RVRT [23], and IART [48].
+Similarly, our findings regarding SCAFFOLD’s performance in highly heterogeneous settings support previous research showing that it is very sensitive to differences in client data [43].
+We observed significant drops and stagnation in our experiments, particularly with the RVRT model, which matches reports that even small changes in local data batches can cause large swings in performance [43].
+SCAFFOLD struggles to train effectively in such diverse scenarios, leading to slower or stalled convergence, which is reflected in our results 3 where its PSNR fluctuates in highly heterogeneous scenarios.
 The increased variance in SCAFFOLD’s performance under skewed data distributions further supports our findings that it fails to adapt smoothly across different client datasets, reinforcing its limitations in federated VSR tasks.
+On top of this, SCAFFOLD requires roughly twice the communication overhead of other algorithms due to its control variates, which further limits its practicality.
+See Section 5.3.6
+Although FedMedian demonstrates better stability than SCAFFOLD in our experiments, it clearly suffers from a lower convergence rate, which is consistent with prior findings.
+Prior work [59] shows that the median-based update is less statistically efficient than mean or FedAvg when each client has a limited local dataset.
+Similarly, it is shown that coordinate-wise median becomes increasingly inefficient as the model grows in size or complexity, making it sub-optimal in high-dimensional settings.
+Additional empirical studies [32] note that FedMedian often converges more slowly because it lacks explicit regularization and tends to produce more conservative updates.
 Overall, these existing studies provide strong theoretical support for our experimental results.
 In contrast, FedVSR’s convergence, stability, and robustness across varying levels of data heterogeneity highlight its effectiveness in overcoming these limitations.
-Qualitative Results:
+5.2.2 Qualitative Results:
 As showen in Fig.
@@
 In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches.
-Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 8.
+5.3 Extended Evaluation and Test Cases
+5.3.1 Additional Baseline: FedSR vs FedVSR
+As discussed in Section 2, existing FL methods for the SISR task were designed for a different setting, providing detailed SR model architectures and introducing extra overhead, which places them in a different category than FedVSR.
+To better understand how our model compares, we tested FedSR with VRT under both heterogeneity settings.
+Since the original paper did not provide a repository, we implemented the code as faithfully as possible.
+The results are shown in Table 3.
+Algorithm Less Heterogeneous More Heterogeneous
+REDS [35] Vid4 [26] REDS [35] Vid4 [26]
+PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
+FedSR [50] 29.70 0.8518 0.1939 24.89 0.7676 0.2470 28.85 0.8297 0.2156 24.70 0.7611 0.2571
+FedVSR 29.96 0.8595 0.1807 24.97 0.7745 0.2284 29.12 0.8371 0.2010 24.78 0.7646 0.2378
+Table 3: Quantitative comparison of FedSR [50], and FedVSR under different heterogeneity settings. Metrics include PSNR, SSIM, and LPIPS for REDS and VID4 datasets.
+5.3.2 Robustness to Communication Failures
+To evaluate the robustness of our method under unreliable communication, we conduct a communication reliability test by simulating scenarios in which a subset of clients fail to upload their updated weights to the server.
+Specifically, we consider failure rates of 25%, 50%, and 75%, meaning that in each round, the corresponding fraction of clients’ updates are discarded.
+We test this setting using two models FedVSR, and FedAvg.
+This setup allows us to compare how each approach handles upload failures and missing updates, and how well they maintain performance under increasingly adverse communication conditions.
+Fig.
+4 illustrates the resulting performance for each method across the different failure rates.
+Our results indicate that FedVSR demonstrates greater resilience, maintaining performance even at high failure rates, whereas FedAvg shows significant degradation as the failure rate increases.
+These findings highlight the advantage of our approach in real-world scenarios where client-server communication may be unreliable.
+Figure 4: FedVSR vs. FedAvg under client upload failures (0–75%), showing FedVSR’s higher robustness.
+5.3.3 Client Population Stress Test
+To evaluate the scalability of our method, we conduct a client population stress test in which we increase the number of clients by three and six times compared to the main experiment 5.1.2, while keeping the client participation rate constant (10%).
+We compare the performance of FedVSR and FedAvg under these conditions.
+This experimental setup allows us to investigate whether our method can maintain high performance when faced with a substantially larger client population, without requiring changes to participation strategies.
+Fig.
+5 shows the results of this test, indicating that FedVSR consistently outperforms FedAvg even as the number of clients increases, demonstrating the scalability and robustness of our approach in large-scale federated learning scenarios.
+Figure 5: FedVSR vs. FedAvg under client population stress test.
+TC PR Algorithm REDS [35] Vid4 [26]
+PSNR SSIM LPIPS PSNR SSIM LPIPS
+Less Heterogeneous
+120 3.3% FedAvg 29.21 0.8261 0.2136 24.36 0.7362 0.2716
+FedVSR 29.36 0.8378 0.1920 24.47 0.7472 0.2431
+240 1.7% FedAvg 28.23 0.8027 0.2327 23.94 0.7079 0.2893
+FedVSR 28.59 0.8155 0.2093 24.10 0.7225 0.2562
+More Heterogeneous
+120 3.3% FedAvg 28.36 0.8028 0.2453 24.09 0.7205 0.2839
+FedVSR 28.54 0.8160 0.2136 24.28 0.7376 0.2531
+240 1.7% FedAvg 26.95 0.7562 0.2979 23.18 0.6641 0.3358
+FedVSR 27.23 0.7742 0.2474 23.44 0.6881 0.2839
+Table 4: Extreme test results for FedAvg and FedVSR under different heterogeneity settings. Here, TC denotes the total number of clients and PR denotes the participation rate.
+5.3.4 Extreme Case: Large Client Population with Low Participation
+To investigate the behavior of our method under extreme conditions, we conduct a test combining a substantially increased client population with very low participation rates.
+Specifically, we evaluate scenarios with three and six times more clients than the main experiment 5.1.2, while reducing the participation rate to approximately one-third (3.3%) and less than one-fifth (1.7%) of the original rate (baseline 10%).
+We compare the performance of FedVSR and FedAvg under these conditions to assess how well each method handles extreme sparsity in client updates combined with a large-scale client population.
+Table 4 presents the results, demonstrating that FedVSR maintains superior performance and robustness even in these challenging scenarios, whereas FedAvg experiences significant performance degradation.
+These findings highlight the resilience and scalability of FedVSR under highly adverse federated learning conditions.
+5.3.5 Effect of Local Epochs on Communication-Efficient Performance
+To investigate the trade-off between local computation and communication efficiency, we compare the performance of FedVSR and FedAvg after 20 global rounds using different numbers of local epochs per client.
+These experiments are measured relative to our main experiment, which consists of 100 global rounds with 1 local epoch per client.
+The results are presented in the Fig.
+6 as percentages of the main experiment’s performance.
+This comparison allows us to assess whether increasing the number of local epochs can compensate for fewer global rounds while highlighting the differences between FedVSR and FedAvg under limited communication.
+Figure 6: Effect of local epochs on FedVSR and FedAvg (% of 100-round FedVSR with 1 local epoch).
+5.3.6 Resource and Efficiency Overhead
+Table 5 reports the computation and memory overhead of different FL algorithms relative to FedAvg.
+We select FedAvg as the baseline because it is a vanilla FL algorithm that introduces no additional overhead.
+To ensure accuracy, each measurement was repeated multiple times.
+We also report results as relative percentages, since our experiments involved three different VSR models with varying computational requirements.
+A value of 0% indicates no overhead beyond FedAvg, while higher values indicate additional computational or memory costs.
+Presenting the results in this way provides a clearer comparison across models.
+For these measurements, we assume an ideal network, as the communication cost of transmitting parameters is negligible compared to training time, as discussed earlier 5.1.5.
+If we were to consider communication cost, in FedVSR each client only needs to send a single scalar, unlike SCAFFOLD or FedSR, which incur a much heavier overhead.
+As shown, FedVSR not only delivers strong performance but also incurs minimal overhead compared to its counterparts.
+   Model    Phase    Latency    Energy    GPU Mem    GPU Util
+   FedVSR    Training    0.25    0.73    0.01    0.17
+   Aggregation    5.39    4.36
+   SCAFFOLD    Training    10.56    11.35    5.53    4.85
+   Aggregation    17.44    21.12
+   FedProx    Training    17.32    19.82    0.71    2.93
+   Aggregation    39.74    39.20
+   FedSR    Training    1.90    4.61    3.16    11.92
+   Aggregation    314.62    620.42
+   FedMedian    Training    0    0    0    0
+   Aggregation    2.15    1.51
+Table 5: Computation and memory overhead of different FL methods relative to FedAvg (%).GPU memory and utilization are reported for both Training and Aggregation phases.
 5.4 Ablation Study
-Evaluating $\mathcal{L}_{\text{HiFr}}$ Loss and Adaptive Aggregation
+5.4.1 Evaluating High Frequency Loss on Centralized Setting
+The results clearly demonstrate that adding $\mathcal{L}_{\text{HiFr}}$ to centralized training negatively impacts performance across all metrics.
+This degradation is observed consistently in both the less and more heterogeneous data splits.
+This outcome confirms that the $\mathcal{L}_{\text{HiFr}}$ loss is not a generic VSR improvement.
+Rather, it is a necessary intervention specifically designed to counteract the core failings of Federated Learning in low-level vision tasks.
+In centralized training, the model has sufficient, high-bandwidth data access to efficiently capture high-frequency details using the standard VSR loss.
+Introducing the DWT-based loss in this setting acts as a redundant constraint, likely leading to over-regularization or gradient noise, which harms the final model quality.
+Our finding establishes that $\mathcal{L}_{\text{HiFr}}$ is a crucial, non-trivial component of the FedVSR framework designed exclusively to fill the performance gap created by the inherent constraints of federated learning for VSR.
+   Algorithm    REDS [35]    Vid4 [26]
+   PSNR    SSIM    LPIPS    PSNR    SSIM    LPIPS
+   Less Heterogeneous
+   VRT    32.20    0.9007    0.1342    26.31    0.8266    0.1780
+[…]
+   More Heterogeneous
+   VRT    31.06    0.8688    0.1505    25.25    0.7900    0.1993
+[…]
+Table 6: Ablation study comparing VRT performance with and without the $\mathcal{L}_{\text{HiFr}}$ loss under centralized training.
+5.4.2 Evaluating High Frequency Loss and Adaptive Aggregation
 To further analyze the impact of the DWT loss term and loss-aware adaptive aggregation, we conducted a series of experiments.
@@
 Similarly, to evaluate loss-aware adaptive aggregation, we first trained VSR models using a greedy aggregation strategy (F-GLA) that prioritizes low-loss updates in all rounds.
-We then compared this approach to an adaptive version that incorporates a decay mechanism (F-ALA).
+We then compared this approach to our adaptive aggregation (F-ALA).
 As shown in Table 7, the DWT loss term consistently outperforms loss-aware adaptive aggregation in all cases.
-Additionally, both methods individually surpass FedAvg [36].
+Additionally, both methods individually surpass FedAvg.
 However, in some cases, the greedy aggregation strategy resulted in only marginal improvements or even a drop in overall performance especially in ”more heterogeneous” setting based on PSNR and SSIM metrics.
 These findings further highlight the importance of high-frequency information not only for improving overall performance but also as a reliable metric for identifying valuable client weights.
-Due to space limitations, LPIPS metric results are provided in the supplementary material 7
 Model Setting Less Heterogeneous More Heterogeneous
-REDS Vid4 REDS Vid4
+REDS [35] Vid4 [26] REDS [35] Vid4 [26]
 PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
@@
 F-GLA 30.57 0.8730 25.42 0.7951 29.50 0.8482 24.99 0.7761
-Table 2: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR and SSIM under different heterogeneity settings.
-Optimal $\lambda$s
-Using a grid search, we explored different values for $\lambda_{\text{VSR}}$ and $\lambda_{\text{HiFr}}$ to determine the optimal ones.
-We varied each parameter from 1 to 0.1 in steps of 0.1 under both bounded and unbounded conditions.
-Due to the high computational cost of the experiments, we trained the VSR models for only 20 rounds and selected the best-performing configurations.
-Our results show that for VRT[30], the optimal values for both parameters are 1 For RVRT[29], the best values are 1 and 0.5, while for IART[54], the optimal values are 0.5 and 0.5, respectively.
-Empirical Fisher for Estimating Weights Importance
-Inspired by [35], we explored estimating the importance of each weight in client models to enhance aggregation using Empirical Fisher (EF) Information.
-EF approximates the Fisher information matrix by leveraging gradients collected during backpropagation after training.
-While this method led to slight performance improvements, it proved computationally too expensive for clients, making it impractical for our setting.
-A more efficient variant of EF was introduced in [52]; however, as it is tailored for classification tasks, we leave its adaptation to VSR for future work.
+Table 7: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation under different heterogeneity settings.
 6 Conclusion
@@
 Despite promising results, a gap remains between FL-based VSR and centralized approaches.
-Future research could refine the aggregation strategy with a more intelligent, task-specific approach tailored for low-level vision.
-Additionally, our loss-guided optimization could be further enhanced to improve stability and better address existing challenges in FL-based VSR.
-Supplementary Material
-7 Extended Evaluation of HiFr Loss and Adaptive Aggregation Using LPIPS
-To complement the analysis based on PSNR and SSIM discussed earlier in the paper, we present here an additional evaluation using LPIPS, which serves as a perceptual similarity metric.
-This acts as an extension to the previously discussed metrics and provides a more complete picture of model performance.
-Due to space limitations, we include this analysis here.
-To evaluate the impact of the DWT loss term and the loss-aware adaptive aggregation strategies, we conducted a set of ablation experiments using FedAvg as the baseline.
-Each component was added independently to isolate its individual effect.
-Specifically, we assessed the contribution of the $\mathcal{L}_{\text{HiFr}}$ term by applying it to FedAvg (F-$\mathcal{L}_{\text{HiFr}}$), and examined two aggregation variants: a greedy loss-aware approach (F-GLA) that prioritizes updates from clients with lower losses, and an adaptive version (F-ALA) that incorporates a decay factor to regulate client influence over time.
-As shown in Table 3 with LPIPS results, the observed trends remain largely consistent across metrics.
-The DWT loss term consistently provides more significant gains compared to the aggregation strategies, and both methods individually outperform the standard FedAvg baseline.
-However, similar to PSNR and SSIM, the greedy strategy sometimes leads to only marginal improvements or even minor performance degradation, especially under high heterogeneity.
-These findings underscore the importance of high-frequency information not only for boosting perceptual quality but also for reliably identifying valuable client updates.
-Model Setting Less Heterogeneous More Heterogeneous
-REDS Vid4 REDS Vid4
-PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
-VRT F-$\mathcal{L}_{\text{HiFr}}$ 29.80 0.8543 0.1910 24.91 0.7681 0.2399 28.98 0.8322 0.2130 24.70 0.7577 0.2500
-F-ALA 29.75 0.8531 0.1912 24.89 0.7681 0.2411 28.90 0.8313 0.2140 24.64 0.7559 0.2520
-F-GLA 29.70 0.8523 0.1914 24.86 0.7670 0.2436 28.83 0.8279 0.2198 24.55 0.7500 0.2555
-RVRT F-$\mathcal{L}_{\text{HiFr}}$ 31.42 0.8888 0.1506 26.02 0.8178 0.1910 30.24 0.8617 0.1612 25.84 0.8042 0.1978
-F-ALA 31.38 0.8880 0.1518 26.00 0.8171 0.1931 30.19 0.8601 0.1728 25.79 0.8041 0.1992
-F-GLA 31.31 0.8879 0.1527 25.95 0.8155 0.1948 29.88 0.8576 0.1807 25.60 0.8032 0.2011
-IART F-$\mathcal{L}_{\text{HiFr}}$ 30.79 0.8801 0.1588 25.46 0.8001 0.1999 29.64 0.8508 0.1899 25.21 0.7837 0.2231
-F-ALA 30.70 0.8789 0.1591 25.44 0.7998 0.2101 29.60 0.8500 0.1903 25.18 0.7831 0.2246
-F-GLA 30.57 0.8730 0.1690 25.42 0.7951 0.2119 29.50 0.8482 0.1929 24.99 0.7761 0.2284
-Table 3: Ablation study on the impact of $\mathcal{L}_{\text{HiFr}}$ and adaptive aggregation on PSNR, SSIM, and predicted LPIPS under different heterogeneity settings.
-8 Further Analysis: Performance Metrics and Visual Comparisons
-This section provides additional performance evaluations for our proposed method.
-Figures 3, 5, and 7 illustrate the PSNR, SSIM and LPIPS trends across different rounds for multiple test sets under various configurations of VRT [30], RVRT [29], and IART [54].
-These results further highlight the effectiveness of our approach in enhancing the performance across different models.
-Figures 4, 6, and 8 provide qualitative visual comparisons on various test frames.
-Each figure presents a reference ground truth (GT) frame alongside low-quality (LQ) input and the outputs generated by different Federated Learning strategies, including FedAvg, FedProx, SCAFFOLD, and our method (FedVSR).
-The reconstructions show that FedVSR produces clearer and more visually faithful outputs, particularly in preserving finer details and reducing artifacts.
-Figure 3: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for VRT.
-Refer to caption (a) Frame 000, Clip 015, REDS
-Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 32.41, 0.8981, 0.1456 Refer to caption (d) FedProx[27] 32.10, 0.8935, 0.1511
-Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 28.85, 0.7625, 0.1934 Refer to caption (g) FedVSR (Ours) 33.79, 0.9169, 0.1151
-Figure 4: Comparison of RVRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
-Figure 5: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for RVRT.
-Refer to caption (a) Frame 040, Clip 1, Vid4
-Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 21.15, 0.7185, 0.3520 Refer to caption (d) FedProx[27] 20.43, 0.6735, 0.4101
-Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 20.41, 0.6679, 0.3803 Refer to caption (g) FedVSR (Ours) 21.19, 0.7200, 0.3433
-Figure 6: Comparison of VRT output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
-Figure 7: PSNR, SSIM and LPIPS values across different rounds for various test sets under different settings for IART.
-Refer to caption (a) Frame 070, Clip 000, REDS
-Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 27.87, 0.8211, 0.2228 Refer to caption (d) FedProx[27] 27.54, 0.8132, 0.2247
-Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 27.69, 0.8151, 0.2207 Refer to caption (g) FedVSR (Ours) 28.07, 0.8319, 0.2014
-Figure 8: Comparison of IART output trained with different Federated Learning algorithms. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
+Future work could improve aggregation with task-specific strategies for low-level vision and enhance our loss-guided optimization for greater stability and effectiveness in FL-based VSR.
+Future Work
+A key direction for future work is extending FedVSR to handle higher-resolution videos and scenarios where clients may contribute data at differing native resolutions.